### v1

# The Origin Symphony: Probing Baryogenesis with Gravitational Waves

Abstract
Affleck-Dine (AD) baryogenesis is compelling yet challenging to probe because of the high energy physics involved. We demonstrate that this mechanism can be realized generically with low-energy new physics without supersymmetry while producing detectable gravitational waves (GWs) sourced by parametric resonance of a light scalar field. In viable benchmark models, the scalar has a mass of ${\cal O}(0.1-10)$ GeV, yielding GWs with peak frequencies of ${\cal O}(10-100)$ Hz. This study further reveals a new complementarity between upcoming LIGO-frequency GW detectors and laboratory searches across multiple frontiers of particle physics.
Introduction.
The origin of the observed matter-antimatter asymmetry $\eta=\frac{n_{b}-n_{\bar{b}}}{n_{\gamma}}\approx 6\times 10^{-10}$ Aghanim et al. (2020) (in terms of baryon-to-photon ratio) remains a long-standing puzzle in cosmology and particle physics, requiring new physics beyond the Standard Model (SM) to resolve it. Among the solutions proposed for the baryon asymmetry puzzle, Affleck-Dine (AD) baryogenesis Affleck and Dine (1985) is a compelling mechanism where the baryon asymmetry originates from the oscillation and subsequent decay of a baryon-number carrying scalar field. In its original form, the AD mechanism is constructed within the framework of supersymmetry (SUSY) Affleck and Dine (1985); Dine et al. (1996); Dine and Kusenko (2003); Enqvist and Mazumdar (2003), where the scalar condensates, potentially essential for baryogenesis, are flat direction moduli fields that are composed of superpartners of the SM quarks/leptons and the Higgs field. These flat direction moduli can generally acquire a large field value displacement during inflation, which spontaneously violates C and CP symmetries, then roll down, oscillate, and decay to baryons in the post-inflationary era. In conventional AD, the scalar field starts to roll and oscillate shortly after the end of inflation, while the inflaton condensate still dominates the Universe before its depletion via reheating Allahverdi and Mazumdar (2012). In alternative models, the AD field itself plays the role of the inflaton Hertzberg and Karouby (2014); Cline et al. (2020a).
Despite the attraction of the AD mechanism, its experimental test is generally challenging due to the high energy scale of the physics involved and the tie to the inflationary era. Fortunately, stochastic gravitational wave background (SGWB) sourced by the AD scalar field provides a new perspective probe by utilizing the naturally high-energy environment in the early cosmos when the AD mechanism operates. As the AD scalar condensate rolls toward its true vacuum and oscillates around it, it can lead to an era of rapid, non-perturbative particle production (e.g., Kitajima et al. (2018); Figueroa and Torrenti (2017); Lozanov and Amin (2019); Hiramatsu et al. (2021); Easther et al. (2007)), known as parametric resonance, and subsequent fragmentation, which can generate SGWB due to the large, time-dependent field inhomogeneities Friedberg et al. (1976); Coleman (1985); Kusenko (1997); Kusenko and Shaposhnikov (1998) produced therein. In the conventional SUSY-based framework, it was shown that parametric resonance of the AD condensates enabling baryogenesis cannot generate detectable GWs Kusenko and Mazumdar (2008); Kusenko et al. (2009); Zhou (2015) 111 In SUSY-based AD models, producing detectable strength of GWs can typically lead to $\eta_{B}\gg 10^{-10}$  Kusenko and Mazumdar (2008); Kusenko et al. (2009). Therefore,  Kusenko and Mazumdar (2008); Kusenko et al. (2009) considered GWs from AD flat-direction fields with $B-L=0$, which is not relevant for baryogenesis. Furthermore, recent lattice simulations revealed that the typical frequency of the GWs generated from SUSY AD models is higher than the currently detectable range Zhou (2015). Finally, the term $|\Phi|^{2}\log(|\Phi|/M_{\mathrm{Pl}})$ in the potential considered in Kusenko and Mazumdar (2008); Kusenko et al. (2009); Zhou (2015) exhibits a pathology at $\Phi=0$ (the impact of which on their GW results warrants re-examination), a pathology absent in our polynomial potential.. On the other hand, Q-balls, generically predicted in SUSY AD models, may result in a period of early matter domination (EMD) that induces GWs at second order through enhancing pre-existing primordial curvature perturbations. White et al. (2021); Kohri and Terada (2018). The Q-ball enhanced GW amplitude strongly depends on the speed of the transition from MD to RD White et al. (2021), with an upper limit given by the commonly used abrupt transition.
The purpose of this work is two-fold: first, we present non-SUSY realizations of the AD mechanism that involve new physics below the electroweak scale and can operate in cosmic epochs long after inflation; second, demonstrate that parametric resonance sourced by AD condensates in such models naturally produces SGWB within reach of upcoming GW experiments. In particular, based on benchmark models with simple renormalizable scalar potentials, we show that in the viable parameter ranges for baryogenesis, parametric resonance of the AD condensate generates GWs with a peak frequency in the range of ${\cal O}(10-100)$ Hz, which can be explored by a combination of upcoming detectors, including Cosmic Explorer (CE) Abbott et al. (2017) and Einstein Telescope (ET) Hild et al. (2008), as well as DECIGO Kawamura et al. (2021) and Big Bang Observatory (BBO) Harry et al. (2006) (to capture the lower frequency tail).
The first discovery of GWs by the LIGO-Virgo-KAGRA collaboration  Abbott et al. (2016a, b); Acernese et al. (2015) ushers in a new era of GW astronomy, followed by the more recent compelling evidence of SGWB identified by pulsar timing array (PTA) observations Carilli and Rawlings (2004); Weltman et al. (2020); Lentati et al. (2015); Babak et al. (2016); Agazie et al. (2023); Afzal et al. (2023); Antoniadis et al. (2023a, b). This study uncovers a generic class of well-motivated sources of SGWB, making a timely complement to the literature. The study also reveals a new potential synergy between GW detection and laboratory experiments at energy, intensity, and neutrino frontiers, such as LHC, DUNE, SHiP, FASER Aad et al. (2021); Aaboud et al. (2018); Bolton et al. (2020); Ariga et al. (2019); Ballett et al. (2020); Abe et al. (2019); Drewes et al. (2018); Barouki et al. (2022): interactions between the AD scalar $\Phi$ and the SM particles are necessary to transfer the asymmetry from the $\Phi$ field to SM baryons, while the characteristic new physics scale in the AD scalar sector is of ${\cal O}(0.1)$${\cal O}(10)$ GeV for viable models with detectable GW frequencies.
The rest of the Letter is organized as follows: in the next section, we present the benchmark models and the prediction for the baryon asymmetry; following that, we show the results of lattice simulation for GW signals from these models. Finally, we conclude with an outlook.
Benchmark Models.
I. AD scalars: background dynamics and asymmetry generation
We consider a complex scalar field $\Phi$, which is a SM gauge singlet with canonical kinetic terms and Einstein gravity. We work with a simple renormalizable polynomial potential:
$V(\Phi)=\lambda_{\Phi}\left|\Phi\right|^{4}+m_{\Phi}^{2}\left|\Phi\right|^{2}-A(\Phi^{n}+\Phi_{\ast}^{n}),$ (1)
with $n$ in principle being any integer $\leq 4$. We choose $n=2$ for concreteness; however, the results are not expected to change significantly with other choices for $n$. The first two terms manifest a $U(1)$ symmetry, which is broken by the third term as necessary to generate the $\Phi$-$\Phi^{\dagger}$ asymmetry, which later transfers to the baryon asymmetry. We consider the general possibility that $\Phi$, as a light spectator field during inflation, starts with a very large (even Planckian) field displacement Starobinsky and Yokoyama (1994); Cui et al. (2024); Hardwick et al. (2017); Markkanen et al. (2018, 2019), e.g., simply by initial condition, or by quantum fluctuation in the inflationary de Sitter space. We have taken into account the isocurvature constraints from CMB data, which is relevant to light spectator fields in general, and found that they can be generally satisfied (see Section .2 of the Supplemental Material for details). We denote the typical value at the end of inflation as $\Phi_{\mathrm{in}}$, the initial value for our calculations  [^2] Other mechanisms can be invoked to drive the $\Phi$ field to large values during inflation, like a non-minimal coupling to gravity or a direct coupling to the inflaton. Our results for the GWs and baryon asymmetry do not depend on the origin of $\Phi_{\mathrm{in}}$.. As long as the (radial) field is light, $H\gg H_{\mathrm{osc}}\simeq\sqrt{V_{,|\Phi|}/|\Phi|}$, where $H_{\mathrm{osc}}$ is the effective mass of $\Phi$ Cui et al. (2024), it remains frozen  [^3] Since we consider small $U(1)$ symmetry-breaking terms, the angular field will always be light when the radial field is light.. The initial phase angle of $\Phi$, $\theta=Arg(\Phi)$, is also set as $\theta_{\mathrm{in}}$ at the end of inflation in our observable patch of the universe. As the universe expands, the Hubble rate drops, and at $H\simeq H_{\mathrm{osc}}$ the field starts rolling. For our parameters, the rolling starts during the radiation-dominated (RD) era, well after the end of inflation.
We analyze the motion of $\Phi$ in terms of its real and imaginary parts (as defined in $\Phi={1\over\sqrt{2}}(\phi_{R}+i\phi_{I})$), which obey the following equations of motion:
$\ddot{\phi}_{{}_{\mathrm{R}}}$ $+3H\dot{\phi}_{{}_{\mathrm{R}}}+m_{{}_{\mathrm{R}}}^{2}\phi_{{}_{\mathrm{R}}}+\lambda_{\Phi}(\phi_{{}_{\mathrm{R}}}^{2}+\phi_{{}_{\mathrm{I}}}^{2})\phi_{{}_{\mathrm{R}}}=0,$ (2)
$\ddot{\phi}_{{}_{\mathrm{I}}}$ $+3H\dot{\phi}_{{}_{\mathrm{I}}}~{}+m_{{}_{\mathrm{I}}}^{2}\phi_{{}_{\mathrm{I}}}~{}+\lambda_{\Phi}(\phi_{{}_{\mathrm{R}}}^{2}+\phi_{{}_{\mathrm{I}}}^{2})\phi_{{}_{\mathrm{I}}}~{}=0,$ (3)
where $m^{2}_{R,I}=m_{\Phi}^{2}\mp 2A$. The Hubble expansion rate in a RD Universe is $H\propto t^{1/2}$. The overdots denote derivatives with respect to cosmic time $t$. We see that the two components have different masses in the presence of $A\neq 0$, and thus, the trajectory in the $\phi_{R}$-$\phi_{I}$ plane is expected to resemble a shrinking ellipse instead of a straight line, even for vanishing initial velocities.
The background field evolution, neglecting the symmetry-breaking term ($A\ll m_{\Phi}^{2}$), is initially dominated by the $\left|\Phi\right|^{4}$ term when the amplitude scales as $\phi\propto 1/a$. When the amplitude drops below $|\Phi_{\ast}|=m_{\Phi}/\sqrt{\lambda_{\Phi}}$ at $t=t_{*}$, the $\left|\Phi\right|^{2}$ term dominates the oscillation and the fields evolve approximately as $\phi_{R,I}\sim a^{-3/2}\cos(m_{R,I}t+c_{R,I})$ (see Section .3 of the Supplemental Material for details). The $\Phi$-$\Phi^{\dagger}$ asymmetry, which will be transferred later to the baryon (or lepton) number, is $n_{\Phi}(t)=i\left(\Phi^{\dagger}\dot{\Phi}-\dot{\Phi}^{\dagger}\Phi\right)=\dot{\phi}_{\mathrm{R}}\phi_{{}_{\mathrm{I}}}-\phi_{{}_{\mathrm{R}}}\dot{\phi}_{\mathrm{I}}$. $n_{\Phi}$ directly corresponds to $n_{B}$ or $n_{L}$ in the absence of significant washout (see discussion later). Using the above approximations for $\phi_{R,I}$ and including the terms $\Gamma_{\Phi}\dot{\phi}_{R,I}$ to model the decay of $\Phi$ to SM states, we can write the co-moving baryon asymmetry generated during the time $t>t_{*}$ as Rubakov and Gorbunov (2017); Lloyd-Stubbs and McDonald (2021, 2023); Mohapatra and Okada (2021):
$\left(\frac{a(t)}{a_{\mathrm{in}}}\right)^{3}n_{B}(t)\simeq 4A\phi_{{}_{\mathrm{R,in}}}\phi_{{}_{\mathrm{I,in}}}\left(\frac{\phi_{\mathrm{in}}}{\phi_{*}}\right)\times$
[…]
where the exponential factor accounts for the decrease in $\Phi$ amplitude due to its decay to SM particles with a constant decay rate $\Gamma_{\Phi}$, enabling the asymmetry transfer to the SM baryons. The specific form of $\Gamma_{\Phi}$ depends on the decay channels, which we will elaborate on later. Evaluating the above integral for $t\rightarrow\infty$ and factoring in the expansion during the RD Universe, we express $n_{B}/s$ as
$\frac{n_{B}}{s}=\begin{cases}\left(\frac{4\alpha^{3}}{\lambda_{\Phi}k_{T_{d}}^{6}}\right)^{1/4}\epsilon_{\Phi}\frac{m_{\Phi}M_{\mathrm{Pl}}}{T_{d}^{2}}\sin(2\theta);&\gamma_{\Phi}\gg 2\epsilon_{\Phi}\\ \left(\frac{\alpha^{3}k_{T_{d}}^{2}}{64\lambda_{\Phi}}\right)^{1/4}\frac{1}{\epsilon_{\Phi}}\frac{T_{d}^{2}}{m_{\Phi}M_{\mathrm{P}}}\sin(2\theta);&\gamma_{\Phi}\ll 2\epsilon_{\Phi}\end{cases}$ (5)
where $M_{\mathrm{Pl}}\simeq 2.4\times 10^{18}$ GeV is the reduced Planck mass, and we define the dimensionless quantities: $\epsilon_{\Phi}=A/m_{\Phi}^{2}$, $\gamma_{\phi}=\Gamma_{\Phi}/m_{\Phi}$ and $k_{T_{d}}=\sqrt{\pi^{2}g(T_{d})/90}$. $T_{d}$ is the temperature of the Universe when the $\Phi$ decay is completed at $H(T_{d})=\Gamma_{\Phi}$, $\alpha$ is the initial fraction of energy density in the spectator sector $\Phi$ relative to the total energy density, and $\theta=\theta_{\mathrm{in}}$ is defined earlier. In case the asymmetry is first transferred to SM leptons, a redistribution factor of $28/79$ due to the EW sphaleron would be added in the RHS of Eq. (5).
Eq. (5) is derived assuming that the asymmetry is generated when the potential is dominated by the $|\Phi|^{2}$ term; thus, we require that the background dynamics prior to $t_{*}$, defined as when $|\Phi(t_{\ast})|=m_{\Phi}/\sqrt{\lambda_{\Phi}}$, is not affected by the $A$ term. The corresponding condition on $A$ can be estimated as follows: The only term in the potential that involves the angular degree of freedom is the symmetry-breaking term $2A|\Phi(t)|^{2}\cos 2\theta(t)\subset V(|\Phi|,\theta)$, indicating an upper bound on the effective mass along the angular direction: $|m^{2}_{\theta}|\leq 2A$. The angular degree of freedom is frozen at $\theta_{\mathrm{in}}$ for $t<t_{*}$, if $|m^{2}_{\theta}|<H^{2}(t_{*})$ Lloyd-Stubbs and McDonald (2023), which translates to $\epsilon_{\Phi}\equiv A/m_{\Phi}^{2}<({\mu^{2}}/{16\alpha})\left({\phi_{\mathrm{in}}}^{2}/{M_{\mathrm{Pl}}}^{2}\right)$, where $\mu^{2}\equiv m_{\Phi}^{2}/(\lambda_{\Phi}\phi_{\mathrm{in}}^{2})$, and $\alpha$ is defined below Eq. 5. As we see in the next section, typically ${\phi_{\mathrm{in}}}\simeq M_{\mathrm{Pl}}$ and $\mu^{2}\lesssim 10^{-5}$ are required for detectable GW signals. With larger $A$, motion along the angular direction is non-negligible, and $\theta(t)$ would have already evolved towards its minimum before $t=t_{*}$, thus reducing $\sin\theta$ and suppressing the asymmetry (see Eq. (5)). For our benchmark examples, the chosen $A$ values satisfy this self-consistency condition.
We now present some numerical examples of the asymmetry. For $\epsilon_{\Phi}=10^{-20}$ the observed $n_{B}/s\sim 10^{-10}$ can be realized with $T_{d}\sim 2.5\times 10^{7}$ GeV when $m_{\Phi}=10\,\mathrm{GeV}$. As we will discuss, transfer of the asymmetry to the baryon sector may involve significant washout (model-dependent), requiring a larger initial $\Phi$ asymmetry to compensate. Therefore we also consider the parameter points that lead to $n_{\Phi}/s\gg 10^{-10}$, for example, with the same $\epsilon_{\Phi}$, but $T_{d}\sim 1$ GeV, one obtains $n_{\Phi}/s\sim 4\times 10^{-5}$. A more detailed parameter scan is given in Section .3 of the Supplemental Material and Fig. S1, demonstrating the possibility of producing a large initial asymmetry to compensate for potential (model-dependent) washout effects.
II. Transfer to SM baryon asymmetry.
The asymmetry in the $\Phi$ field generated during its oscillation needs to eventually be transferred to the SM baryon asymmetry via its subsequent decay to baryons or leptons (further assisted by electroweak sphalerons in the latter case). This can generally be realized by a B- or L-violating coupling involving $\Phi$ in terms of effective higher dimensional operators, such as $\Phi QQQL/\Lambda^{3}$ Allahverdi and Mazumdar (2012), $\Phi UDDUDD/\Lambda^{6}$ and $\Phi LLELLE/\Lambda^{6}$. Whether the related constraints, for instance, from $n-\bar{n}$ oscillation, can be satisfied depends on the specifics of the UV completion, including flavor structure in the couplings. As a proof of principle, in the following, we outline examples of renormalizable UV complete models realizing the $\Phi$ to baryon asymmetry transfer and summarize the conditions on viable model parameters. Further details of the analysis are given in Section .4 of the Supplemental Material. The core results of this work are independent of the details of the asymmetry transfer mechanism.
(A) $\Phi$ decay to leptons (leptogenesis).
A simple renormalizable model involves:
$\mathcal{L}\supset y_{N,~{}i}\Phi\bar{N_{i}}^{C}{N_{i}}+g_{\nu,~{}ij}\bar{N}_{i}HL_{j}+{\mathrm{h.c.}},$ (6)
where $N_{i}$’s are right-handed (RH) Majorana neutrinos (with $i=1,2,3$ as a typical choice), $H$ is SM Higgs, $L_{j}$ are SM lepton doublets. This model potentially relates to neutrino mass generation, which nevertheless is not imposed as a requirement here. We consider the simple scenario where $\Phi$ decays to $NN$, then $N$ transfers the asymmetry to SM leptons via freeze-out or freeze-in processes enabled by the $NHL$ coupling, as considered in the literature of leptogenesis Akhmedov et al. (1998); Klarić et al. (2021a, b); Cui et al. (2011); Hambye and Teresi (2016); Flood et al. (2022)  [^4] Unlike in leptogenesis, here we do not require additional CP violation or out-of-equilibrium condition in neutrino sector to generate the asymmetry, as the asymmetry is already generated in $\Phi$..
(B) $\Phi$ decay to baryons (direct baryogenesis)
A simple renormalizable model, in this case, includes:
$\mathcal{L}\supset y_{\chi}\Phi\chi\chi+g_{i}\chi\bar{u}_{i}\varphi+\lambda_{ij}\varphi d_{i}d_{j}+{\mathrm{h.c.}},$ (7)
where $\chi$ is a singlet Majorana fermion, $\varphi$ is an up-type diquark scalar, and the flavor indices $i,~{}j=1,2,3$ in down-type quark combination $d_{i}d_{j}$ should be antisymmetric. In the viable parameter region, the decay proceeds with $\Phi\rightarrow\chi\chi$, followed by $\chi\rightarrow udd$.
Both models are subject to cosmological conditions which impose lower limits on $m_{\Phi}$ and $m_{\chi}$ or $m_{N}$. For $\Phi$ to decay to baryons (before BBN), it is subject to kinematic conditions: $m_{\Phi}\gtrsim 2$ GeV, and $m_{\chi}\gtrsim 1$ GeV. For $\Phi$ asymmetry to be transferred to the SM leptons via $N$ before the EW phase transition, $m_{\Phi}>2m_{N}\gtrsim O(0.1)$ GeV Klarić et al. (2021a). In addition, for both models, the potential asymmetry washout effects should be either suppressed or can be compensated by a larger initial $\Phi$ asymmetry produced during its oscillation (which has been shown viable in the last section). Furthermore, relevant laboratory experiments at energy and intensity frontiers searching for B-violation, color-charged new particles, sterile neutrinos, etc., may provide additional constraints on model parameters. In Section .4 of the Supplemental Material, we elaborate on relevant constraints and demonstrate examples of viable parameter regions. On the other hand, as we will see in our GW results, an upper limit in mass of $m_{\Phi}\lesssim O(10)$ GeV corresponds to the currently detectable GW peak frequency range, which is below $\sim O(100)$ Hz. This, combined with the aforementioned cosmological lower limit of $m_{\Phi}\gtrsim O(0.1-1)$ GeV, demonstrates that intriguingly the parameter range of our interest points to new physics at the energy scale of $O(0.1-10)$ GeV, indicating a new potential complementarity between GW detection and laboratory searches across particle physics frontiers.
Parametric Resonance and GW Signals. We have seen that the generation of the baryon asymmetry is determined by the classical motion of $\Phi$ as an averaged background field. In addition to this classical motion of $\Phi$, we must consider the corresponding fluctuations around it, which leads to parametric resonance that may source GWs. We proceed by decomposing $\phi_{R,I}=\bar{\phi}_{R,I}(t)+\delta\phi_{R,I}(x,t)$, where $\bar{\phi}_{R,I}$ describe the classical background motion and $\delta\phi_{R,I}(x,t)$ correspond to the fluctuations, which have a quantum origin. The evolution of background fields $\bar{\phi}_{R,I}$ follows Eqs. (2), (3), whereas we expand the fluctuations in Fourier modes as
$\delta\phi_{R}(x,t)=\int{d^{3}k\over(2\pi)^{3}}\left[a_{R,k}e^{i\vec{k}\cdot\vec{x}}\delta\phi_{R,k}(t)+h.c.\right]$ (8)
and similarly for $\delta\phi_{I}(x,t)$.
As the background fields oscillate around the minimum of the potential, the fluctuations in momentum-space evolve according to
$\ddot{\delta\phi}_{R,I}+3H\dot{\delta\phi}_{R,I}+m_{R,I}^{2}{\delta\phi_{R,I}}+{k^{2}\over a^{2}}{\delta\phi_{R,I}}$
$+\lambda_{\phi}(3\bar{\phi}_{R,I}^{2}+\bar{\phi}_{I,R}^{2})\delta\phi_{R,I}+2\bar{\phi}_{I}\bar{\phi}_{R}\delta\phi_{I,R}=0$ (9)
We thus see that, in general, the equations for $\delta\phi_{R,I}$ are coupled. We first consider the representative limit of $A=0$, where $m_{R}=m_{I}=m_{\Phi}$ and the background field trajectory is simply a straight line through the origin (for vanishing initial velocities, as is the case here).
In this case, we can rotate our field basis and define fields in the radial direction (along the background motion) ϕsubscriptitalic-$\phi_{\parallel}$ and perpendicular to it (orthogonal direction) as ϕsubscriptitalic-$\phi_{\perp}$. By construction, $\bar{\phi}_{\perp}=0$, and the fluctuation equations for ϕsubscriptitalic-$\phi_{\parallel}$ and ϕsubscriptitalic-$\phi_{\perp}$ decouple, namely,
$\ddot{\delta\phi}_{\parallel,\perp}+3H\dot{\delta\phi}_{\parallel,\perp}+\left({k^{2}\over a^{2}}+m_{\Phi}^{2}+q_{\parallel,\perp}\lambda_{\Phi}\bar{\phi}_{\parallel,\perp}^{2}\right)\delta\phi_{\parallel,\perp}=0$ (10)
where $q_{\parallel,\perp}=\{3,1\}$. As the background field oscillates quasi-periodically around the origin, fluctuations can be exponentially amplified for certain ranges of wavenumbers. Through the exponential amplification, the occupation number of these modes grows rapidly, and they can thus be treated classically Polarski and Starobinsky (1992); Kofman et al. (1997). This phenomenon is referred to as “preheating” Greene et al. (1997) in the context of primordial particle production following inflation, while it also can occur for oscillating spectator scalar fields in general (see, for instance, our related work Cui et al. (2024)).
To quantify parametric resonance, we start by defining a dimensionless field amplitude $\tilde{\phi}=\phi/\phi_{\mathrm{in}}$ and a corresponding conformal scaling for time as $\mathrm{d}\eta=\sqrt{\lambda_{\Phi}}\phi_{\mathrm{in}}\mathrm{d}t/a$ with $\phi_{\mathrm{in}}$ being the initial amplitude of the scalar field $\phi_{\mathrm{in}}=|\Phi|_{\mathrm{in}}$. The dimensionless fluctuation equation becomes (derivatives are over $\eta$)
${\delta\phi}_{\parallel,\perp}^{\prime\prime}+2\tilde{H}{\delta\phi}_{\parallel,\perp}^{\prime}+\left({\tilde{k}^{2}}+a^{2}\mu^{2}+q_{\parallel,\perp}a^{2}\tilde{\phi}^{2}\right)\delta\phi_{\parallel,\perp}=0$ (11)
where $\tilde{k}^{2}=k^{2}/\lambda\phi_{\mathrm{in}}^{2}$ and $\tilde{\phi}=\bar{\phi}/\phi_{\mathrm{in}}$.
[…]
The growth of fluctuations and their back-reaction break the homogeneous oscillation of $\Phi$, leading to the inhomogeneous structure of scalar fields with an effective anisotropic stress tensor $\Pi_{ij}^{TT}=\nabla_{i}\phi^{a}\nabla_{j}\phi^{a}$ (where the repeated index `$a$’ indicates the number of the scalar components and a summation is implied). These time-dependent field inhomogeneities source (sub-Hubble) gravitational waves whose frequency depends on the characteristic energy scale involved $H_{c}$. The peak amplitude and frequency of GW spectra generated from sources peaked at some characteristic scale $k_{c}$ in momentum-space can be estimated as Giblin and Thrane (2014)
$f_{\mathrm{GW}}^{\mathrm{peak}}=$ $2.7\times 10^{10}\sqrt{\frac{H_{c}}{M_{\mathrm{P}}}}\frac{k_{c}}{H_{c}}\,{\mathrm{Hz}}\,,$ (12)
$\Omega_{\mathrm{GW}}^{\mathrm{peak}}=$ $2.3\times 10^{-4}\,\alpha^{2}\,\beta\,w\left(\frac{k_{c}}{\sigma}\right)\left(\frac{H_{c}}{k_{c}}\right)^{2}.$ (13)
where $\alpha$ is a measure of the fraction of the energy in the GW source here the scalar field inhomogeneities and this $\alpha$ coincides with the $\alpha$ we defined below Eq. 5 relative to the total energy density of the Universe, and $\beta$ describes the degree of anisotropy of the source which must be determined from simulations. In typical scenarios $\beta=\mathcal{O}(0.01-0.1)$. The equation of the state of the Universe at the time of the production of GWs for a radiation-dominated Universe is $w=1/3$ and $\sigma$ is the width of the source in momentum-space, which can be taken as $\sigma\sim k_{c}$ for peaked sources (see Ref. Giblin and Thrane (2014) for the derivation and e.g. Refs. Cui and Sfakianakis (2023); Cui et al. (2024) for further applications of Eqs (12) and (13)).
The dynamics of the generated GWs are governed by the linearized equation for the transverse-traceless tensor perturbation sourced by the above anisotropic stress
$\ddot{h}_{ij}+3H\dot{h}_{ij}-\frac{\nabla^{2}h_{ij}}{a^{2}}=\frac{2}{a^{2}M_{\mathrm{P}}^{2}}\Pi_{ij}^{TT},$ (14)
The quantity of interest is the energy density power spectrum of the GWs normalized by the critical energy density of the Universe today. For GWs produced in a radiation-dominated era, the spectrum of the energy density of GWs (per logarithmic momentum interval) observable today is Easther and Lim (2006); Easther et al. (2007, 2008); Dufaux et al. (2007); Cui et al. (2024); Garcia-Bellido and Figueroa (2007); Garcia-Bellido et al. (2008):
$\Omega_{\mathrm{G}W,0}h^{2}$ $=\frac{h^{2}}{\rho_{\mathrm{crit}}}\frac{d\rho_{\mathrm{G}W}}{d\ln k}\Bigg{|}_{t=t_{0}}=\frac{h^{2}}{\rho_{\mathrm{crit}}}\frac{d\rho_{\mathrm{G}W}}{d\ln k}\Bigg{|}_{t=t_{\mathrm{e}}}\frac{a_{\mathrm{e}}^{4}\rho_{\mathrm{e}}}{a_{0}^{4}\rho_{\mathrm{crit},0}}$
$=\Omega_{\mathrm{rad},0}h^{2}\Omega_{\mathrm{GW},\mathrm{e}}\left(\frac{g_{\ast}}{g_{0}}\right)^{-1/3},$ (15)
where the critical density of the Universe is $\rho_{\mathrm{crit}}=3M_{\mathrm{P}}^{2}H^{2}$ and $\Omega_{\mathrm{rad},0}h^{2}=h^{2}\rho_{\mathrm{rad},0}/\rho_{\mathrm{crit}}=4.3\times 10^{-5}$. In particular, we employ the publicly available code Cosmolattice Figueroa et al. (2021, 2023) in a three-dimensional lattice, in order to numerically solve the metric perturbation equations in Eq. (14) along with the equations for the scalar system in an expanding universe.
The benchmark examples of observed spectra today from our model simulation are shown in Fig 1. For the initially $\Phi^{4}$ dominant system, as in our case, the field starts rolling/oscillating when $H\sim\sqrt{\lambda_{\Phi}}\phi_{\mathrm{in}}$. This combination of $\sqrt{\lambda_{\Phi}}\phi_{\mathrm{in}}$ also determines the frequency of the observed GWs, according to Eq. 12. We consider the RD universe with the scalar field system contributing to a fraction of the total energy $\alpha$ and $\alpha=0.3$ is assumed in Fig 1. According to Eq. 13, $\Omega^{\mathrm{peak}}_{\mathrm{GW}}\propto\alpha^{2}\propto\phi^{8}_{\mathrm{in}}$, and we found that a sizable $\alpha\sim O(0.1)$ which optimizes GW amplitude requires $\phi_{\mathrm{in}}\sim M_{\mathrm{P}}$. As we have seen from the discussion on the resonance structure of the model, the parameters $\mu_{\pm}$ (or simply $\mu$ when $A=0$) determine the nature and efficiency of resonance Greene et al. (1997). We take $\mu=10^{-5}$ as an example for stable resonance, which fixes $\lambda_{\Phi}\phi_{\mathrm{in}}^{2}$. with the optimal choice of $\phi_{\mathrm{in}}\sim M_{\mathrm{Pl}}$, we find $\lambda_{\Phi}\sim 10^{-30}$ for the mass range of interest. As discussed earlier, $m_{\Phi}\gtrsim O(0.1)$ GeV is required for $\Phi$ to successfully transfer the asymmetry to SM baryons. Interestingly, in the viable range of $m_{\Phi}\sim O(0.1)-O(10)$ GeV, the resulting peak GW frequency lies well within the range of ET/CE and with sufficient amplitudes, making signal detection promising for the near future. A higher range of $m_{\Phi}$ is compatible with baryogenesis but requires futuristic higher frequency GW detectors Aggarwal et al. (2021) to capture the peak frequency range of the signal. Due to the finite box size and lattice spacing, low-frequency modes well off the peak cannot be resolved in our simulation, which is a common issue in the lattice simulation literature Easther et al. (2007). Therefore, to extend the spectral prediction to lower frequencies, we apply the standard assumption that at low wave numbers, the GW spectrum grows as $k^{3}$, as derived from causality arguments Easther et al. (2007); Caprini et al. (2009); Hook et al. (2021). This is shown in dashed lines in Fig. 1. Depending on $m_{\Phi}$, the low-$k$ tail may be detectable by DECIGO and BBO, providing the prospect for a correlated detection across multiple frequency bands, probing the peak and tail of the spectrum simultaneously.
Figure 1: We show the GW spectrum originating from the AD model considered here for three benchmark masses of $\Phi$. The potential parameters $\lambda_{\Phi}=10^{-30}$, $A=0$ are chosen to ensure efficient parametric resonance (see main text for details). The fraction of initial scalar field energy density over the total energy density of the Universe is $\alpha=30\%$. The dashed lines at lower frequencies are extrapolated according to the causal super-horizon $k^{3}$ scaling.
Finally, note that since $\Phi$ is coupled to SM fermions in order to transfer the asymmetry to baryons, the effect of parametric excitation of these fermionic states should be evaluated. Nevertheless, in the context of preheating, it was found that during scalar field oscillations, the energy transfer to bosons is typically faster and larger than to fermions due to Pauli blocking Greene and Kofman (1999); Giudice et al. (1999); Greene and Kofman (2000); Sorbo (2001); Adshead and Sfakianakis (2015). Therefore, we expect our resonance analysis to remain robust in the presence of fermion couplings.
Conclusion and Outlook. In this Letter, we first demonstrate that the Affleck-Dine mechanism for baryogenesis can be realized by a complex light scalar field $\Phi$ with a mass $m_{\Phi}$ well below the electroweak scale without invoking supersymmetry. In particular, we use a convex, renormalizable, polynomial potential (which does not lead to the formation of Q-balls generally present in SUSY AD models). Such a scalar field acts as a spectator field during inflation while acquiring a large field value at the end of inflation. The subsequent rolling down and oscillation of the scalar field occur long after inflation, during the radiation-dominated era, triggering the generation of baryon asymmetry. Parametric resonance and fragmentation generally arise during the oscillation of the scalar field, naturally leading to GW production. We demonstrate that the theory may lead to successful baryogenesis while sourcing detectable GW signals with a peak frequency of O(10-100) Hz, within reach of experiments such as ET and CE. Furthermore, the characteristic new physics scale, characterized by $m_{\Phi}$, intriguingly lies in the range of ${\cal O}(0.1\mathrm{-}10)$ GeV, while the transfer of the $\Phi$ asymmetry to the SM B- or L-asymmetry requires interactions between the new physics sector and the SM states. Hence, from this well-motivated scenario, a new, natural complementarity arises between SGWB detection and laboratory searches for new particle physics across the energy, intensity, and neutrino frontiers. The specifics of the complementary laboratory signal depend on the details of the asymmetry transfer mechanism. Examples include: searches for ${\cal O}(0.1\mathrm{-}10)$ GeV mass sterile neutrino with DUNE, SHiP, FASER, etc. and exotic multi-jet signals at the LHC (see Section .4 of the Supplemental Material). This work thus sheds new light on probing early Universe baryogenesis through its GW imprint, potentially complemented by associated laboratory probes, while presenting a well-motivated new physics source for SGWB.
Acknowledgements. We thank Brian Shuve for helpful discussions. YC is supported by the US Department of Energy under award number DE-SC0008541. AG is supported by ”Excellence Initiative - Research University (2020-2026)” IDUB grant. PS is supported by Grant-in-Aid for Scientific Research (B) under Contract No. 23K25873 (23H01177). Numerical computation in this work was carried out at the Yukawa Institute Computer Facility. EIS acknowledges support by the U.S. Department of Energy, Office of Science, Office of High Energy Physics program under Award Number DE-SC0022021.
Supplemental Material
Yanou Cui, Anish Ghoshal, Pankaj Saha and Evangelos I. Sfakianakis
.1 Inflationary fluctuations and initial conditions
In some versions of the Affleck-Dine mechanism, an extra coupling is introduced in the potential in order to generate a radial minimum at large field values Allahverdi and Mazumdar (2012). We argue that this is not necessary for the AD field to acquire a large initial value. It has been known that a light scalar field in de-Sitter space-time acquires quantum fluctuations, eventually leading to an equilibrium distribution Starobinsky and Yokoyama (1994). In the case of a quartic real scalar field with $V={1\over 4}\lambda\phi^{4}$, the variance of the equilibrium distribution is $\langle\phi^{2}\rangle\sim H^{2}/\sqrt{\lambda}$, whereas for a quadratic field $V={1\over 2}m^{2}\phi^{2}$ we arrive at $\langle\phi^{2}\rangle\sim H^{4}/m^{2}$. As an example, $\langle\phi^{2}\rangle\sim M_{\mathrm{Pl}}^{2}$ and $0.1\,\mathrm{GeV}\lesssim m\lesssim 10\,\mathrm{GeV}$ requires an inflationary Hubble scale of at least $10^{-10}M_{\mathrm{Pl}}\lesssim H\lesssim 10^{-9}M_{\mathrm{Pl}}$ for quadratic fields. If the quartic term dominates during inflation with $\lambda\sim 10^{-30}$ a Planckian standard deviation requires $H\sim 10^{-15}M_{\mathrm{Pl}}$ or greater. The variance of the distribution grows as $\langle\phi^{2}\rangle=(H^{2}/4\pi^{2})N$ where $N$ is the $e$-folding number. We thus see that an equilibrium distribution with a Planckian standard deviation requires much more than the usual minimum amount of $60$ $e$-folds of inflation. This may be considered as a way to infer whether a prolonged inflationary phase took place in the early universe.
The above discussion (see Ref. Cui et al. (2024) for more details) applies to the absolute value of the complex scalar field $|\phi|$. The argument will enter its equilibrium distribution much sooner, and the distribution itself will be much simpler: uniform between 00 and $2\pi$. The intricacies of multi-dimensional random walks and their applications to de Sitter fluctuations of fields with continuous symmetries (including complex scalars) can be found in Ref. Adshead et al. (2020). Finally, a large initial field value $\phi_{\mathrm{in}}$ could simply result from an ad hoc initial condition. One may argue that the initial condition for the inflaton itself is set ad hoc, although quantum fluctuations can be invoked to explain it. Most importantly, our results are independent of the details of how $\phi_{\mathrm{in}}$ is realized (the same is true for the inflationary attractor).
.2 Isocurvature perturbations
In our previous work Cui et al. (2024), we briefly discussed the possible isocurvature fluctuations from spectator scalar fields during inflation. In the current case, there are two degrees of freedom due to the complex nature of the AD field. If we decompose the complex scalar into a radial and an angular degree of freedom, the radial field will have similar behavior to a real scalar (for slight quantitative differences, see Adshead et al. (2020)), with isocurvature fluctuations that scale as ${\cal P}_{\cal S}\sim H_{I}^{2}/M_{\mathrm{Pl}}^{2}$ if we take the initial value of the radial field to be ${\cal O}(M_{\mathrm{Pl}})$ originating solely form de-Sitter fluctuations as described in Section .1. We thus see that for reasonable values of the inflationary Hubble scale $H_{I}$, these fluctuations are suppressed. Furthermore, if the AD field decays to SM particles, we can expect the isocurvature perturbation to vanish.
However, the angular degree of freedom, which is responsible for the baryon number, also receives de Sitter fluctuations during inflation. This means that the AD mechanism proceeds differently in different parts of the Universe due to the fluctuations in the initial condition for the angular field. This leads to baryon isocurvature fluctuations, which are more constrained by the measurement of the CMB.
We follow previous computations of isocurvature fluctuations in models of AD baryogenesis, in particular, Ref. Hertzberg and Karouby (2014); see also Kasuya et al. (2008). During inflation we can decompose the AD field $\Phi$ into a radial component $R$ and an angular component $\theta$. We consider a typical value of the radial field $\bar{R}=\sqrt{\langle R^{2}\rangle}\sim M_{\mathrm{P}l}$ and thus
$\langle\delta\theta^{2}\rangle\simeq{1\over 4\pi^{2}}{H_{I}^{2}\over\bar{R}^{2}}$ (S1)
Ref. Hertzberg and Karouby (2014) introduced a multiplicative ${\cal O}(1)$ factor to account for the two-dimensional random walk of a complex scalar (see also Ref. Adshead et al. (2020)). Since the baryon number is proportional to $\sin(n\theta)$ the normalized first order fluctuations $\delta n_{b}$ follow
${\delta n_{b}\over n_{b}}=n\delta\theta\cot(\theta)$ (S2)
The corresponding power of the baryon isocurvature temperature fluctuations is
$\left\langle\left(\delta T\over T\right)\right\rangle_{\mathrm{isoc}}\simeq{0.004}{\Omega_{b}^{2}\over\Omega_{m}^{2}}{n^{2}H^{2}\over\bar{R}^{2}}\cot^{2}(n\theta)$ (S3)
We can take $\cot^{2}(n\theta)\sim 1$ and $\Omega_{b}/\Omega_{m}\simeq 0.16$, leading to
$\left\langle\left(\delta T\over T\right)\right\rangle_{\mathrm{isoc}}\simeq 10^{-4}{n^{2}H^{2}\over\bar{R}^{2}}$ (S4)
Taking a simple model of single-field inflation for simplicity and following again Ref. Hertzberg and Karouby (2014)
$\left\langle\left(\delta T\over T\right)\right\rangle_{\mathrm{adiabatic}}\simeq{1\over 20}{H^{2}\over 8\pi^{2}M_{\mathrm{Pl}^{2}}\epsilon}={1\over 20}{H^{2}\over 8\pi^{2}M_{\mathrm{Pl}^{2}}\epsilon}\simeq 4\times 10^{-5}{H^{2}\over M_{\mathrm{P}l}^{2}}{1\over r}$ (S5)
where we used the tensor-to-scalar ratio rather than the first slow roll parameter, using $r=16\epsilon$. The ratio of the power in this isocurvature over adiabatic perturbations is then
$\alpha_{II}\sim n^{2}r$ (S6)
where we dropped ${\cal O}(1)$ terms and used $\bar{R}\sim M_{\mathrm{Pl}}^{2}$. We see that the Planck bound $\alpha_{II}<3.9\times 10^{-2}$ is satisfied for $n=2$ and $r<10^{-2}$, which is, for example satisfied for Starobinsky inflation, Higgs(-like) inflation and some $\alpha$-attractor models, where $r={\cal O}(10^{-3})$. As the inflationary scale is lowered, this bound is more easily satisfied.
.3 Estimate of the asymmetry
We provide a more detailed derivation for the asymmetry calculation than the one found in the main text, following Refs. Lloyd-Stubbs and McDonald (2021, 2023). The scalar potential for the complex field $V(\Phi)=\lambda_{{}_{\Phi}}\left|\Phi\right|^{4}+m_{{}_{\Phi}}^{2}\left|\Phi\right|^{2}-A(\Phi^{2}+\Phi_{\ast}^{2})$ with quadratic symmetry-breaking terms is
$V(\Phi)=\lambda_{{}_{\Phi}}\left|\Phi\right|^{4}+m_{{}_{\Phi}}^{2}\left|\Phi\right|^{2}-A(\Phi^{2}+\Phi_{\ast}^{2})$ (S7)
The field equations for the real and imaginary parts are:
$\ddot{\phi}_{{}_{\mathrm{R}}}$ $+3H\dot{\phi}_{{}_{\mathrm{R}}}+m_{{}_{\mathrm{R}}}^{2}\phi_{{}_{\mathrm{R}}}+\lambda_{{}_{\Phi}}(\phi_{{}_{\mathrm{R}}}^{2}+\phi_{{}_{\mathrm{I}}}^{2})\phi_{{}_{\mathrm{R}}}=0,$ (S8)
$\ddot{\phi}_{{}_{\mathrm{I}}}$ $+3H\dot{\phi}_{{}_{\mathrm{I}}}~{}+m_{{}_{\mathrm{I}}}^{2}\phi_{{}_{\mathrm{I}}}~{}+\lambda_{{}_{\Phi}}(\phi_{{}_{\mathrm{R}}}^{2}+\phi_{{}_{\mathrm{I}}}^{2})\phi_{{}_{\mathrm{I}}}~{}=0,$ (S9)
where
$m_{{}_{\mathrm{R}}}^{2}=m_{{}_{\Phi}}^{2}-2A;\quad m_{{}_{\mathrm{I}}}^{2}=m_{{}_{\Phi}}^{2}+2A.$ (S10)
while the Hubble expansion is that of a radiation-dominated Universe ($H\propto t^{1/2}$.)
At large field values (defined as $|\Phi|>\phi^{*}=m_{\Phi}/\sqrt{\lambda_{\Phi}}$), the symmetry-breaking term is subleading, and we consider the dominant term of the potential
$V(\Phi)\simeq\lambda_{\Phi}|\Phi|^{4}\,,~{}~{}~{}|\Phi|>\phi_{*},$ (S11)
while at small field values, the quartic term is negligible, and we may approximate the potential as
$V(\Phi)\simeq m_{\Phi}^{2}|\Phi|^{2}-A\left(\Phi^{2}+(\Phi^{\ast})^{2}\right)\,,~{}~{}~{}|\Phi|<\phi_{*}$ (S12)
By defining time $t_{*}$ as $|\Phi(t_{*})|=\phi_{*}$, we approximate the field evolution during the two eras as Turner (1983):
When $t<t_{\ast}$:
$\phi_{\mathrm{R}}(t<t_{\ast})$ $=$ $\left(\frac{a_{\mathrm{i}n}}{a}\right)\phi_{\mathrm{R},in}$ (S13)
$\phi_{\mathrm{I}}(t<t_{\ast})$ $=$ $\left(\frac{a_{\mathrm{i}n}}{a}\right)\phi_{\mathrm{I},in}$ (S14)
$\phi_{{}_{\mathrm{R}}}(t>t_{\ast})$ $=$ $\phi_{{}_{\mathrm{R,\ast}}}\left(\frac{a_{\ast}}{a}\right)^{3/2}\cos(m_{{}_{\mathrm{R}}}(t-t_{\ast})),$ (S15)
$\phi_{{}_{\mathrm{I}}}(t>t_{\ast})$ $=$ $\phi_{{}_{\mathrm{I,\ast}}}~{}\left(\frac{a_{\ast}}{a}\right)^{3/2}\cos(m_{{}_{\mathrm{I}}}(t-t_{\ast}))$ (S16)
where $\phi_{{}_{\mathrm{R,\ast}}}=(a_{\mathrm{in}}/a_{*})\phi_{\mathrm{R},in}$ and $\phi_{{}_{\mathrm{I,\ast}}}=(a_{\mathrm{in}}/a_{*})\phi_{\mathrm{I},in}$. The conserved charge density in the $\Phi$ condensate is
$n_{0}(t)=i\left(\Phi^{\dagger}\dot{\Phi}-\dot{\Phi}^{\dagger}\Phi\right)=\dot{\phi}_{\mathrm{R}}\phi_{{}_{\mathrm{I}}}-\phi_{{}_{\mathrm{R}}}\dot{\phi}_{\mathrm{I}}$ (S17)
where we have used the subscript ‘00’ to denote the generated asymmetry in the $\Phi$ sector alone, without considering the decay to SM particles.
Substituting the approximate solutions for the background fields, the asymmetry becomes
$n_{0}(t)=\phi_{{}_{\mathrm{R,\ast}}}\phi_{{}_{\mathrm{I,\ast}}}\left(\frac{a_{\ast}}{a}\right)^{3}\left[m_{{}_{\mathrm{I}}}\sin(m_{{}_{\mathrm{I}}}(t-t_{\ast}))\cos(m_{{}_{\mathrm{R}}}(t-t_{\ast}))-m_{{}_{\mathrm{R}}}\sin(m_{{}_{\mathrm{R}}}(t-t_{\ast}))\cos(m_{{}_{\mathrm{I}}}(t-t_{\ast}))\right]\,.$ (S18)
Since the symmetry-breaking term is subdominant to the other terms in the potential, we consider $A\ll m_{{}_{\Phi}}^{2}$ and expand the asymmetry $n_{0}(t)$ at leading order in $A/m_{{}_{\Phi}}^{2}$
$n_{0}(t)=m_{\Phi}\phi_{{}_{\mathrm{R,\ast}}}\phi_{{}_{\mathrm{I,\ast}}}\left(\frac{a_{\ast}}{a}\right)^{3}\left[\sin\left(\frac{2A(t-t_{\ast})}{m_{{}_{\Phi}}}\right)+\frac{A}{m_{{}_{\Phi}}^{2}}\sin(2m_{{}_{\Phi}}(t-t_{\ast}))\right]\,.$ (S19)
The second term is highly oscillatory and will amount to zero when averaged over several field oscillations. Hence, for $t>t_{\ast}$ the expression in Eq. (S19) reduces to
$n_{0}(t)=\phi_{{}_{\mathrm{R,\ast}}}\phi_{{}_{\mathrm{I,\ast}}}\left(\frac{\phi_{{}_{\mathrm{in}}}}{\phi_{{}_{\ast}}}\right)\left(\frac{a_{\mathrm{i}n}}{a}\right)^{3}m_{{}_{\Phi}}\sin\left(\frac{2A(t-t_{\ast})}{m_{{}_{\Phi}}}\right)\,.$ (S20)
The generated asymmetry will oscillate around zero with a period $T_{\mathrm{asy}}=\pi m_{{}_{\Phi}}/A$. Defining the comoving asymmetry $n_{{}_{C}}(t)\equiv(a(t)/a_{\mathrm{i}n})^{3}n(t)$, we can estimate the transfer of generated asymmetry to a conserved SM asymmetry via $B$-conserving $\Phi$ decay to SM particles with a constant rate $\Gamma_{\Phi}$.
Finally, the transferred asymmetry to the SM can be calculated as
$\tilde{n}_{{}_{C}}(t)=\int_{t_{\ast}}^{t}{\mathrm{d}}t~{}\Gamma_{\Phi}n_{{}_{\mathrm{C}}}(t)=\frac{\Gamma_{\Phi}\phi_{{}_{\ast}}^{2}\sin(2\theta)m_{{}_{\Phi}}}{2}\int_{t_{\ast}}^{t}{\mathrm{d}}t~{}e^{-\Gamma_{\Phi}(t-t_{\ast})}\sin\left(\frac{2A(t-t_{\ast})}{m_{{}_{\Phi}}}\right)$ (S21)
with $\phi_{{}_{\mathrm{R,\ast}}}\phi_{{}_{\mathrm{I,\ast}}}=\phi_{{}_{\ast}}^{2}\sin\theta\cos\theta=\frac{1}{2}\phi_{{}_{\ast}}^{2}\sin 2\theta$. The final expression for the transferred asymmetry for $t-t_{*}\gg m_{\Phi}/A$ is
$\tilde{n}_{{}_{C}}=\frac{\Gamma_{\Phi}\phi_{{}_{\ast}}^{2}\sin(2\theta)m_{{}_{\Phi}}^{2}}{2A}\left(1+\left(\frac{\Gamma_{\Phi}m_{{}_{\Phi}}}{2A}\right)^{2}\right)^{-1}.$ (S22)
This expression has two limiting cases depending on whether the $\Phi$ lifetime is short ($\Gamma_{\Phi}\gg 2A/m_{{}_{\Phi}}$) or long ($\Gamma_{\Phi}\ll 2A/m_{{}_{\Phi}}$) compared to the period of oscillation of the asymmetry:
$\tilde{n}_{{}_{C}}=\begin{cases}\frac{2A}{\Gamma_{\Phi}}\,\phi_{{}_{\ast}}^{2}\sin(2\theta)\quad&\Gamma_{\Phi}\gg 2A/m_{{}_{\Phi}}\\ \frac{\Gamma_{\Phi}m_{{}_{\Phi}}^{2}}{2A}\,\phi_{{}_{\ast}}^{2}\sin(2\theta)\quad&\Gamma_{\Phi}\ll 2A/m_{{}_{\Phi}}\end{cases}\,.$ (S23)
In the absence of efficient washout (as discussed in the main text and Supplemental Material Sect. .4), $\Phi$ asymmetry $\tilde{n}_{{}_{C}}/s$ corresponds directly to baryon or lepton number asymmetry, which we assume below for simplicity. In case of a large washout effect, $\tilde{n}_{{}_{C}}/s$ would be larger than the resultant $n_{B}/s$. We do not elaborate on how the washout affects the prediction in $n_{B}/s$ in the latter case, as it is highly dependent on the details of the asymmetry transfer mechanism. Instead, we will demonstrate the feasibility of generating $\tilde{n}_{{}_{C}}/s\gtrsim 10^{-10}$ that can compensate for a potentially large washout suppression.
Ultimately, the quantity $n_{B}/s$ needs to match the observed value
$\frac{n_{B}}{s}\Bigg{|}_{\mathrm{obs}}=0.861\pm 0.005\times 10^{-10}.$
To this end, we multiply the comoving transferred asymmetry with $(a_{\ast}/a)^{3}$ to compute the total asymmetry
$\tilde{n}=\left(\frac{a_{\ast}}{a}\right)^{3}\tilde{n}_{{}_{C}}$ (S24)
For the RD universe, as in our case, the equation of state is $w=1/3$ leading to
$\left(\frac{a_{\ast}}{a}\right)^{3}=\left(\frac{H(t)}{H_{\ast}}\right)^{3/2}$ (S25)
Dividing $\tilde{n}$ by the entropy density $s=4k_{T}^{2}T^{3}$, with $k_{T}=\sqrt{\pi^{2}g(T)/90}$, we arrive at the baryon-to-photon ratio
$\frac{n_{B}}{s}=\begin{cases}\left(\frac{4\alpha^{3}}{k_{T_{d}}^{6}\lambda_{{}_{\Phi}}}\right)^{1/4}\left(\frac{A}{m_{{}_{\Phi}}^{2}}\right)\left(\frac{m_{{}_{\Phi}}M_{\mathrm{P}}}{T_{d}^{2}}\right)\sin(2\theta)\quad&\Gamma_{\Phi}\gg 2A/m_{{}_{\Phi}}\\ \left(\frac{\alpha^{3}k_{T_{d}}^{2}}{64\lambda_{{}_{\Phi}}}\right)^{1/4}\left(\frac{m_{{}_{\Phi}}^{2}}{A}\right)\left(\frac{T_{d}^{2}}{m_{{}_{\Phi}}M_{\mathrm{P}}}\right)\sin(2\theta)\quad&\Gamma_{\Phi}\ll 2A/m_{{}_{\Phi}}\,,\end{cases}$ (S26)
where the $\Phi$ decay completes when $\Gamma_{\Phi}=H(T_{d})=k_{T_{d}}T_{d}^{2}/M_{\mathrm{P}}$ and the Hubble scale at $t_{*}$ is $H_{\ast}\simeq\frac{3}{4}{m_{{}_{\Phi}}^{2}}/{\lambda_{{}_{\Phi}}}$. We can also elucidate the dependence on $\phi_{\mathrm{in}}$, by rewriting the above expression as
$\frac{n_{B}}{s}=\begin{cases}\mu\left(\frac{4\lambda_{{}_{\Phi}}\alpha^{3}}{k_{T_{d}}^{6}}\right)^{1/4}\left(\frac{A}{m_{{}_{\Phi}}^{2}}\right)\left(\frac{\phi_{\mathrm{in}}M_{\mathrm{P}}}{T_{d}^{2}}\right)\sin(2\theta)\quad&\Gamma_{\Phi}\gg 2A/m_{{}_{\Phi}}\\ \mu\left(\frac{\alpha^{3}k_{T_{d}}^{2}}{64\lambda_{{}_{\Phi}}^{3}}\right)^{1/4}\left(\frac{m_{{}_{\Phi}}^{2}}{A}\right)\left(\frac{T_{d}^{2}}{\phi_{\mathrm{in}}M_{\mathrm{P}}}\right)\sin(2\theta)\quad&\Gamma_{\Phi}\ll 2A/m_{{}_{\Phi}}\,,\end{cases}$ (S27)
where for sufficient resonance, we require $\mu^{2}\equiv m_{{}_{\Phi}}^{2}/(\lambda_{{}_{\Phi}}\phi_{\mathrm{in}}^{2})\lesssim 10^{-5}$.
In the case where the asymmetry in the $\Phi$ field is first transferred to a SM lepton asymmetry $n_{L}$, which is later converted to a baryon asymmetry $n_{B}$ via the sphaleron process, the above result gets multiplied by a factor of $28/79$ since Harvey and Turner (1990); Asaka et al. (2002); Garcia-Bellido et al. (1999)
$\frac{n_{B}}{s}=-\frac{28}{79}\frac{n_{L}}{s}\,.$ (S28)
Figure S1 shows the resulting asymmetry as a function of the temperature of the universe at the time of complete $\Phi$ decay for different values of the potential parameters; $m_{\Phi}$ and $A$. The quartic term is chosen to keep $m_{\Phi}^{2}/\lambda_{\Phi}\phi_{\mathrm{in}}^{2}=10^{-5}$ for sufficient resonance. We see the strong dependence of $n_{B}/s$ on $A$ and $T_{d}$, as expected. This allows the model to be viable, even in cases where significant washout occurs, by increasing the initially produced asymmetry through proper parameter choices.
Finally, let us reiterate one assumption that was made in the above analytic results for the generated asymmetry. As mentioned in the main text, our analysis assumes that the motion of $\Phi$ is radial at early times, for $\lambda_{\Phi}|\Phi|^{4}\gg m_{\Phi}^{2}|\Phi|^{2}$. After that, the potential is dominated by the quadratic terms, which in our case includes the bare mass and the asymmetry $A(\Phi^{2}+\Phi_{*}^{2})$. By simply equating the quartic and bare mass terms, we arrive at the time of transition between the two regimes of $t_{*}$ when $|\Phi(t_{*})|=m_{\Phi}/\sqrt{\lambda_{\Phi}}$. Demanding that at early times (before $t_{*}$) the asymmetry term can be neglected leads to $\epsilon_{\Phi}=A/m_{\phi}^{2}<(\mu^{2}/15\alpha)(\phi_{\mathrm{in}}^{2}/M_{\mathrm{Pl}}^{2})$ where $\mu^{2}\equiv m_{\Phi}^{2}/(\lambda_{\Phi}\phi_{\mathrm{in}}^{2})\lesssim 10^{-5}$. For $\phi_{\mathrm{in}}\sim M_{\mathrm{Pl}}$ the above inequality can be written as $\epsilon_{\Phi}\lesssim 10^{-5}$. If we push $\epsilon_{\Phi}$ above this value, the model is not invalided, but the asymmetry would get an extra suppression factor since $\sin\theta<\sin\theta_{\mathrm{in}}$. The amount of suppression depends on the amount of non-radial motion for $t<t_{*}$. That being said, the computed $n_{B}/s$ can be many orders of magnitude above the measured value of $10^{-10}$, meaning that a suppression of $\sin\theta$ is not “fatal” for the model, even for $\epsilon_{\Phi}>10^{-5}$. We do not perform a detailed analysis for this regime, which simply involves computing the angular motion of $\Phi$ from the first instant when it starts rolling for a given set of potential parameters.
As an example, for the benchmark values of $\lambda_{\Phi}\sim 10^{-30}$ and $m_{\Phi}=10\,\mathrm{GeV}$, we must have $\epsilon_{\Phi}\lesssim 10^{-6}$ for the validity of our estimates in Eq. (5). The temperature at the decay can vary between $T_{d}^{\mathrm{max}}\sim 10^{8}\,\mathrm{GeV}$ for instantaneous decay and $T_{d}^{\mathrm{min}}=10\mathrm{MeV}$, if the decay occurs right before BBN. Therefore, to obtain the correct ratio of $n_{B}/s$, we require $\epsilon_{\Phi}=10^{-38}(T_{d}/\mathrm{GeV})^{2}$ when $\gamma_{\Phi}\ll\epsilon_{\Phi}$.
Figure S1: The maximum baryon-to-photon ratio $\left.{n_{B}/s}\right|_{\mathrm{max}}={n_{\Phi}/s}$ (neglecting possible washout and the sphaleron factor) for three different masses of the scalar field as in Fig (1) as a function of the temperature $T_{d}$ of the $\Phi$ decay. The solid and dashed lines correspond to $\epsilon_{\Phi}=10^{-20}$ and $\epsilon_{\Phi}=10^{-6}$ respectively. The maximum temperature (denoted as the dots) for each line corresponds to the case of instantaneous decay of the scalar at $t=t_{\ast}$. The horizontal dot-dashed line corresponds to the observed value of the baryon asymmetry.
.4 Transfer of the $\Phi$ asymmetry to the SM baryon asymmetry
In this section we elaborate the example models realizing the transfer of $\Phi$ asymmetry to the SM baryon asymmetry and the related constraints.
(A) $\Phi$ decay to leptons (leptogenesis).
A simple renormalizable model involves:
$\mathcal{L}\supset y_{N,~{}i}\Phi\bar{N_{i}}^{C}{N_{i}}+g_{\nu,~{}ij}\bar{N}_{i}HL_{j}+{\mathrm{h}.c.}$ (S29)
where $N_{i}$’s are right-handed (RH) Majorana right-handed neutrinos, $H$ is SM Higgs, $L_{j}$ are SM lepton doublets. The number of $N$ species, $N_{N}$, is a model-dependent parameter, and one can take $N=1$ for simplicity or $N_{N}=3$ as an example by analogy to the SM neutrinos. Again, while it is appealing to simultaneously address the origin of neutrino masses, it is not imposed as a requirement for our consideration.
We consider the simple scenario where $\Phi$ decays to $NN$, then $N$ transfers the asymmetry to SM leptons via freeze-out or freeze-in processes as considered in leptogenesis literature Akhmedov et al. (1998); Klarić et al. (2021a, b); Cui et al. (2011); Flood et al. (2022); Shuve and Yavin (2014); Hambye and Teresi (2016). As seen from our GW results, $m_{\Phi}\lesssim O(10)$ GeV is preferred in order to yield detectable GW frequency below $\sim 100$ Hz. This implies that light $N$ of mass $\lesssim O(1-10)$ GeV. In this mass range, conventional $N$ decays to SM leptons via a 3-body or loop-suppressed process, which typically occurs after electroweak sphaleron already turns off, albeit before BBN. The washout effect from inverse decay or back-scattering is also significant in this low mass range. However, as found in recent leptogenesis literature, asymmetry transfer from $N$ to SM leptons can be realized while sphaleron is still active by freeze-out (generalized out-of-equilibrium decay or scattering) or freeze-in (via $N$-$\nu_{\mathrm{S}M}$ mixing/oscillation) mechanisms, for $m_{N}$ as low as O(100) MeV. The strength of the washout effect can be highly flavor-dependent. Even in the presence of large washout suppression, as shown in the main text, our Affleck-Dine model can initially generate a larger $\Phi$ asymmetry, which compensates for the washout to eventually result in the observed baryon asymmetry. Unlike with the leptogeneis case, here, the $NHL$ interaction only serves the purpose of transferring the $\Phi$ ($N$) asymmetry to SM leptons instead of generating the asymmetry via CP violation in the SM neutrino sector. Therefore, the details of the asymmetry evolution in our case may differ from its leptogenesis counterparts and are highly model-dependent, which may be elucidated by dedicated studies beyond the scope of this work. Furthermore, an alternative to the Majorana $N$ as shown in Eq. S29 is to introduce quasi-Dirac $N$ where the washout effect can be more suppressed for GeV mass range (also see Cline et al. (2020b) for this possibility and the related phenomenological implications). For all of the above possibilities for $N$ to SM asymmetry transfer, $\Phi\rightarrow NN$ decay is required to occur before EW sphaleron turns off, i.e. $\Gamma_{\Phi\rightarrow NN}\sim y_{\Phi}^{2}m_{\Phi}/(16\pi)\gtrsim H(T_{\mathrm{E}W})$. This leads to the constraint $y_{\Phi}\gtrsim 10^{-7}\sqrt{\frac{\mathrm{GeV}}{m_{\Phi}}}$, which can be easily satisfied.
Laboratory signatures of the case of $\Phi$ decaying to leptons generally relate to sterile neutrino searches in the mass range of $O(0.1-10)$ GeV, which are being conducted at a range of experiments such as DUNE, SHiP, FASER et al. ste (2022); Bolton et al. (2020); Ariga et al. (2019); Ballett et al. (2020); Abe et al. (2019); Drewes et al. (2018); Barouki et al. (2022).
(B) $\Phi$ decay to baryons.
A simple renormalizable model in this case includes:
$\mathcal{L}\supset y_{\chi}\phi_{\chi}\chi\chi+g_{i}\chi\bar{u}_{i}\varphi+\lambda_{ij}\varphi d_{i}d_{j}+{\mathrm{h}.c.}$ (S30)
where $\chi$ is a singlet Majorana fermion, $\varphi$ is an up-type diquark scalar, and the flavor indices $i,~{}j=1,2,3$ in down-type quark combination $d_{i}d_{j}$ should be antisymmetric. This type of model finds its close analogy in supersymmetric theories, in particular, $\chi$ may relate to neutralino or singlino Cui and Sundrum (2013), $\varphi$ an up-type squark, $\lambda_{ij}$ as UDD type of R-parity violating (RPV) coupling. Therefore, existing searches for RPV SUSY may lead to constraints on masses and couplings that we consider here. The LHC constraints on the colored particle $\phi$ depend on whether it resembles the first two generations of squark or stop, and depends on whether RPV decay $\phi\rightarrow dd$ or R-parity conserving decay $\phi\rightarrow u\chi$ dominates (assuming $\chi$ reveals itself as MET) Karmakar (2023); Sirunyan et al. (2018, 2021); Aad et al. (2021). In any case, generally, $m_{\varphi}\gtrsim$1 TeV is the ballpark of the current LHC constraints, which we will use for our estimates. This constraint determines that in the parameter range of our interest where $m_{\chi}\sim O(1-10)$ GeV, $\chi$ decays proceeds with 3-body channel $\chi\rightarrow udd$. For simplicity, we assume $g_{i}\sim 1$. B-violating coupling $\lambda_{ij}$ is subject to various constraints by laboratory searches at the intensity frontier. This model does not involve L-violation and does not induce proton decay, which usually introduces constraints on B-violating couplings Arnold et al. (2013). Due to the antisymmetry in $i,~{}j$, $n$-$\bar{n}$ constraint is rather weak, while di-nucleon decay $pp\rightarrow K^{+}K^{+}$ provides the strongest bound Goity and Sher (1995); Barbier et al. (2005) of $\lambda_{12}\lesssim 10^{-6}$, for $m_{\varphi}\sim$ TeV, $m_{\chi}\sim$ O(GeV). However, the bound is relaxed for heavy flavor quarks, e.g. $\lambda_{23}\lesssim 1$.
The kinematic condition of $\chi$ decaying to baryons leads to lower limit on masses: $m_{\Phi}>2$ GeV, $m_{\chi}>1$ GeV. The decays of both $\Phi\rightarrow\chi\chi$ and $\chi\rightarrow udd$ are required to occur before BBN, i.e., $\Gamma_{\Phi\rightarrow\chi\chi}\sim\frac{1}{16\pi}y_{\chi}^{2}m_{\Phi}>H(T_{\mathrm{BBN}})$, and $\Gamma_{\chi\rightarrow udd}>H(T_{\mathrm{BBN}})$. We then estimate the constraint on coupling parameters as: $y_{\chi}\gtrsim 10^{-12}\frac{\mathrm{G}eV}{m_{\phi}}$, $\sum_{i,j}g_{i}\lambda_{ij}\gtrsim\sqrt{10^{-21}\frac{m^{4}_{\varphi}\mathrm{GeV}}{m_{\chi}^{5}}}$. With $m_{\varphi}\sim$ TeV, $m_{\chi}\sim$ O(GeV), we find that $\sum_{i,j}g_{i}\lambda_{ij}\gtrsim 10^{-5}$. If $\lambda_{12}$ dominates $\chi$ decay, this would contradict the aforementioned di-nucleon decay bound. Therefore, we choose to consider the pattern of heavy flavor domination, which is well motivated and has been widely considered in SUSY context Brust et al. (2012a, b); Cui et al. (2011); Franceschini and Torre (2013). In particular, we consider the dominant decay of $\chi$ involves b-quark $\chi\rightarrow udb$ or $\chi\rightarrow usb$. As $m_{b}\simeq 4$ GeV, the kinematics requires $m_{\chi}>$4 GeV and $m_{\Phi}>$8 GeV. The involvement of heavy b-quark also alleviates the potential washout suppression of baryon asymmetry due to inverse decay $udd\rightarrow\chi$ and back-scattering $udd\rightarrow\bar{u}\bar{d}\bar{d}$: by requiring $\chi$ decay below $T\sim m_{b}$ (compatible with the condition of decay before BBN), the washout effect gets Boltzmann suppression and quickly becomes inefficient.
Laboratory signatures of the case of $\Phi$ decaying to baryons depend on model details. One interesting possibility is a multi-jet signal at the LHC (eight jets with b-tagging), potentially with displaced vertices, from pair-production of $\phi$ followed by $\phi\rightarrow\chi u\rightarrow uudd$, which is distinct yet relates to existing searches such as Aaboud et al. (2018).
.5 Asymmetry dependence of GW production
It is worth exploring in more detail the connection between the asymmetry of the potential and the shape and amplitude of the resulting GW spectrum. Let us start from the symmetric case $A=0$, where the potential becomes
$V(\phi_{R},\phi_{I})=\frac{1}{4}\lambda_{\Phi}(\phi_{R}^{2}+\phi_{I}^{2})^{2}+\frac{1}{2}m_{\Phi}^{2}(\phi_{R}^{2}+\phi_{I}^{2})=\frac{1}{4}\lambda_{\Phi}\phi_{R}^{4}+\frac{1}{4}\lambda_{\Phi}\phi_{I}^{4}+\frac{1}{2}m_{\Phi}^{2}\phi_{R}^{2}+\frac{1}{2}m_{\Phi}^{2}\phi_{I}^{2}+\frac{1}{2}\lambda_{\Phi}\phi_{R}^{2}\phi_{I}^{2}$ (S31)
Since the potential is radially symmetric, we can redefine the fields as $\phi_{\parallel}=\phi_{R}\cos\theta-\phi_{I}\sin\theta$ and $\phi_{\perp}=\phi_{R}\sin\theta+\phi_{I}\cos\theta$, where at the background level $\bar{\phi}_{\perp}=0$. This essentially makes the analysis follow our earlier work Cui et al. (2024), where we considered (among others) systems with two scalar fields with a quartic coupling, where one field is initially displaced from its minimum, and the other is (classically) at zero. This specific example is a combination of Models A and D of Ref. Cui et al. (2024). The equations of motion for the fluctuations in the $\phi_{\parallel}-\phi_{\perp}$ basis are
Figure S2: The spectra of the scalar field fluctuations along (left) and perpendicular (right) to the direction of the background trajectory in the case of a rotationally symmetric potential with $A=0$, $\lambda_{\Phi}=5\times 10^{-35}$ and $m_{\Phi}=0.1$ GeV.
$\ddot{\delta\phi_{\parallel}}+3H\dot{\delta\phi_{\parallel}}+\left({k^{2}\over a^{2}}+m_{\Phi}^{2}+3\lambda_{\Phi}\bar{\phi}^{2}\right)\delta\phi_{\parallel}=0$ (S32)
$\ddot{\delta\phi_{\perp}}+3H\dot{\delta\phi_{\perp}}+\left({k^{2}\over a^{2}}+m_{\Phi}^{2}+\lambda_{\Phi}\bar{\phi}^{2}\right)\delta\phi_{\perp}=0$ (S33)
In principle, we can see two sources of parametric resonance: self-resonance of the ϕsubscriptitalic-$\phi_{\parallel}$ field due to the quartic term $3\lambda_{\Phi}\bar{\phi}^{2}\delta\phi_{\parallel}$ and resonant amplification of the ϕsubscriptitalic-$\phi_{\perp}$ field through the coupling term $\lambda_{\Phi}\bar{\phi}^{2}\delta\phi_{\perp}$.
The background field trajectory is (by construction) along $\bar{\phi}_{\perp}=0$ and thus a straight line through the origin on the $\phi_{R}-\phi_{I}$ plane. Floquet analysis of a quartic field in a nonexpanding universe predicts that the resonance is stronger for $q=1$  ($\mu_{k}^{max}\sim 0.15$) than $q=3$ ($\mu_{k}^{max}\sim 0.04$Greene et al. (1997), where $q$ is defined as the coefficient before $\lambda_{\Phi}\bar{\phi}^{2}$ in Eqs. S33. Figure S2 shows the spectra of $\delta\phi_{\parallel},\delta\phi_{\perp}$ for different times, where we see that the $\delta\phi_{\perp}$ spectrum dominates, as expected from the standard analysis of preheating of quartic fields Greene et al. (1997). In the $\phi_{R}-\phi_{I}$ basis, the spectra will be identical for $\theta=\pi/4$, since each spectrum has an equal contribution from $\delta\phi_{\parallel}$ and $\delta\phi_{\perp}$. If the initial angle between the real and imaginary parts of $\Phi$ changes, the spectra of $\delta\phi_{\parallel}$ and $\delta\phi_{\perp}$ remain unchanged, whereas the spectra of $\phi_{R}$ and $\phi_{I}$ change, since the inner product between the basis vectors of the two bases change with the angle $\theta$. That being said, the resulting GW spectrum is unchanged, as expected, due to the rotational symmetry of the potential.
We now move to the case of $A\neq 0$. We parametrize the effect by $\epsilon_{\Phi}=A/m_{\Phi}^{2}$. As expected, small values of the angle do not significantly affect the background trajectory and the resonance structure. As seen in Fig. S3, for $\epsilon_{\phi}\leq 0.005$, the GW spectrum is independent of the initial angle $\theta_{\mathrm{in}}$ and actually indistinguishable from the symmetric case with $\epsilon_{\Phi}=0$. This means that for the asymmetry values that we use for our benchmark examples $\epsilon_{\Phi}<10^{-5}$, the resonance is “blind” with respect to the exact value of asymmetry in the potential.
Figure S3: The resulting GW spectra for different values of the asymmetry parameter $A$ (color-coded according to the legend) and initial angle $\theta\equiv\arctan(\phi_{I}/\phi_{R})|_{\mathrm{init}}=\pi/6,\pi/4,\pi/3$ (solid, dashed and dotted respectively).
That being said, it is worth pushing the asymmetry value higher in order to cover the entire possible parameter space. For $0.05\lesssim\epsilon_{\Phi}\lesssim 0.15$ we see an enhancement of the GW spectrum at large values of the frequency $f\gtrsim 100$ Hz, while the independence of $\theta_{\mathrm{in}}$ still holds. This is an intriguing observation, as it provides an additional motivation for GW detectors that can probe frequencies around $1$ kHz with a sensitivity similar to CE and ET Zhang et al. (2023); Chen et al. (2023). For $\epsilon_{\Phi}\simeq 0.25$ the UV enhancement of the GW spectrum remains, but we also see a difference between the initial angle choices between the real and imaginary parts of $\Phi$. This means that a possible detection of a spectrum with such a UV enhancement would suffer from a partial degeneracy between $\epsilon_{\phi}$ and $\theta_{\mathrm{in}}$. Since we focused on $\epsilon_{\Phi}\ll 1$, such a highly model-dependent calculation is outside the scope of the present work. Finally, for $\epsilon_{\Phi}\gtrsim 0.3$, parametric resonance is severely suppressed, rendering the GW signals below the detection threshold.

### v2

@@
-The Origin Symphony: Probing Baryogenesis with Gravitational Waves
+Probing baryogenesis with gravitational waves
 Abstract
-Affleck-Dine (AD) baryogenesis is compelling yet challenging to probe because of the high energy physics involved.
-We demonstrate that this mechanism can be realized generically with low-energy new physics without supersymmetry while producing detectable gravitational waves (GWs) sourced by parametric resonance of a light scalar field.
-In viable benchmark models, the scalar has a mass of ${\cal O}(0.1-10)$ GeV, yielding GWs with peak frequencies of ${\cal O}(10-100)$ Hz.
-This study further reveals a new complementarity between upcoming LIGO-frequency GW detectors and laboratory searches across multiple frontiers of particle physics.
+Affleck-Dine (AD) baryogenesis is compelling yet challenging to probe because of the high-energy physics involved.
+We demonstrate that this mechanism can be generically realized with low-energy new physics without supersymmetry while producing detectable gravitational waves (GWs) sourced by parametric resonance of a light scalar field.
+In viable models, the scalar has a mass of $\mathcal{O}(0.1-10)$ GeV, yielding GWs with peak frequencies of $\mathcal{O}(10-100)$ Hz.
+This study further reveals a new complementarity between upcoming LIGO-frequency GW detectors and laboratory searches across frontiers of particle physics.
 Introduction.
-The origin of the observed matter-antimatter asymmetry $\eta=\frac{n_{b}-n_{\bar{b}}}{n_{\gamma}}\approx 6\times 10^{-10}$ Aghanim et al.
-(2020) (in terms of baryon-to-photon ratio) remains a long-standing puzzle in cosmology and particle physics, requiring new physics beyond the Standard Model (SM) to resolve it.
-Among the solutions proposed for the baryon asymmetry puzzle, Affleck-Dine (AD) baryogenesis Affleck and Dine (1985) is a compelling mechanism where the baryon asymmetry originates from the oscillation and subsequent decay of a baryon-number carrying scalar field.
-In its original form, the AD mechanism is constructed within the framework of supersymmetry (SUSY) Affleck and Dine (1985); Dine et al.
-(1996); Dine and Kusenko (2003); Enqvist and Mazumdar (2003), where the scalar condensates, potentially essential for baryogenesis, are flat direction moduli fields that are composed of superpartners of the SM quarks/leptons and the Higgs field.
-These flat direction moduli can generally acquire a large field value displacement during inflation, which spontaneously violates C and CP symmetries, then roll down, oscillate, and decay to baryons in the post-inflationary era.
-In conventional AD, the scalar field starts to roll and oscillate shortly after the end of inflation, while the inflaton condensate still dominates the Universe before its depletion via reheating Allahverdi and Mazumdar (2012).
-In alternative models, the AD field itself plays the role of the inflaton Hertzberg and Karouby (2014); Cline et al.
-(2020a).
-Despite the attraction of the AD mechanism, its experimental test is generally challenging due to the high energy scale of the physics involved and the tie to the inflationary era.
-Fortunately, stochastic gravitational wave background (SGWB) sourced by the AD scalar field provides a new perspective probe by utilizing the naturally high-energy environment in the early cosmos when the AD mechanism operates.
-As the AD scalar condensate rolls toward its true vacuum and oscillates around it, it can lead to an era of rapid, non-perturbative particle production (e.g., Kitajima et al.
+The observed excess of matter over antimatter in the universe, in terms of the baryon-to-photon ratio $\eta=\frac{n_{b}-n_{\bar{b}}}{n_{\gamma}}\approx 6\times 10^{-10}$ Aghanim et al.
+(2020), remains a long-standing puzzle in particle physics and cosmology.
+Affleck-Dine (AD) baryogenesis Affleck and Dine (1985) is a compelling solution to the puzzle, where the asymmetry emerges from the oscillation and decay of a baryon number-carrying complex scalar field.
+Originally formulated with supersymmetry (SUSY) Affleck and Dine (1985); Dine et al.
+(1996); Dine and Kusenko (2003); Enqvist and Mazumdar (2003), it involves flat direction moduli fields composed of superpartners of standard model (SM) particles.
+These light moduli fields can acquire large displacements during inflation, break C and $CP$ symmetries and subsequently roll down and oscillate to produce baryon asymmetry, while the inflaton still dominates the Universe Allahverdi and Mazumdar (2012) (alternatively, the AD field itself can be the inflaton Hertzberg and Karouby (2014); Cline et al.
+(2020a)).
+Despite its appeal, experimentally testing the AD mechanism is challenging because of the high-scale physics involved in the inflationary era.
+Fortunately, the stochastic gravitational wave background (SGWB) generated by the oscillating AD field offers a promising new probe by utilizing the naturally high-energy environment in the early cosmos: oscillating scalar fields can trigger parametric resonance, i.e., rapid nonperturbative particle production, leading to large inhomogeneities and fragmentation that source GWs (e.g., Kitajima et al.
 (2018); Figueroa and Torrenti (2017); Lozanov and Amin (2019); Hiramatsu et al.
 (2021); Easther et al.
-(2007)), known as parametric resonance, and subsequent fragmentation, which can generate SGWB due to the large, time-dependent field inhomogeneities Friedberg et al.
-(1976); Coleman (1985); Kusenko (1997); Kusenko and Shaposhnikov (1998) produced therein.
-In the conventional SUSY-based framework, it was shown that parametric resonance of the AD condensates enabling baryogenesis cannot generate detectable GWs Kusenko and Mazumdar (2008); Kusenko et al.
-(2009); Zhou (2015) 111 In SUSY-based AD models, producing detectable strength of GWs can typically lead to $\eta_{B}\gg 10^{-10}$  Kusenko and Mazumdar (2008); Kusenko et al.
+(2007); Friedberg et al.
+(1976); Coleman (1985); Kusenko (1997); Kusenko and Shaposhnikov (1998)).
+In conventional SUSY AD baryogenesis, parametric resonance does not generate detectable GWs Kusenko and Mazumdar (2008); Kusenko et al.
+(2009); Zhou (2015)   [^1] In SUSY AD models, producing detectable GWs can typically lead to $\eta_{B}\gg 10^{-10}$ Kusenko and Mazumdar (2008); Kusenko et al.
 (2009).
 Therefore, Kusenko and Mazumdar (2008); Kusenko et al.
-(2009) considered GWs from AD flat-direction fields with $B-L=0$, which is not relevant for baryogenesis.
-Furthermore, recent lattice simulations revealed that the typical frequency of the GWs generated from SUSY AD models is higher than the currently detectable range Zhou (2015).
-Finally, the term $|\Phi|^{2}\log(|\Phi|/M_{\mathrm{Pl}})$ in the potential considered in Kusenko and Mazumdar (2008); Kusenko et al.
-(2009); Zhou (2015) exhibits a pathology at $\Phi=0$ (the impact of which on their GW results warrants re-examination), a pathology absent in our polynomial potential..
-On the other hand, Q-balls, generically predicted in SUSY AD models, may result in a period of early matter domination (EMD) that induces GWs at second order through enhancing pre-existing primordial curvature perturbations.
-White et al.
-(2021); Kohri and Terada (2018).
-The Q-ball enhanced GW amplitude strongly depends on the speed of the transition from MD to RD White et al.
-(2021), with an upper limit given by the commonly used abrupt transition.
-The purpose of this work is two-fold: first, we present non-SUSY realizations of the AD mechanism that involve new physics below the electroweak scale and can operate in cosmic epochs long after inflation; second, demonstrate that parametric resonance sourced by AD condensates in such models naturally produces SGWB within reach of upcoming GW experiments.
-In particular, based on benchmark models with simple renormalizable scalar potentials, we show that in the viable parameter ranges for baryogenesis, parametric resonance of the AD condensate generates GWs with a peak frequency in the range of ${\cal O}(10-100)$ Hz, which can be explored by a combination of upcoming detectors, including Cosmic Explorer (CE) Abbott et al.
-(2017) and Einstein Telescope (ET) Hild et al.
-(2008), as well as DECIGO Kawamura et al.
-(2021) and Big Bang Observatory (BBO) Harry et al.
-(2006) (to capture the lower frequency tail).
-The first discovery of GWs by the LIGO-Virgo-KAGRA collaboration  Abbott et al.
+(2009) considered GWs from AD flat-direction fields with BL=0B\!-\!L=0, not relevant for baryogenesis.
+Furthermore, the potential term $|\Phi|^{2}\log(|\Phi|/M_{\mathrm{Pl}})$ considered in Kusenko and Mazumdar (2008); Kusenko et al.
+(2009); Zhou (2015) exhibits a pathology at $\Phi=0$, which is absent in our polynomial potential; its impact on GW emission warrants re-examination”.
+Meanwhile, SUSY AD models can lead to Q-balls, which may result in early matter domination and induce GWs at second order by enhancing preexisting curvature perturbations Kusenko and Mazumdar (2008); Kusenko et al.
+(2009).
+Q-ball-induced GWs strongly depend on the speed of the transition from matter to radiation domination White et al.
+(2021).
+The purpose of this work is twofold: first, we present non-SUSY realizations of the AD mechanism that involve new physics below the electroweak scale and can operate in cosmic epochs long after inflation; second, we demonstrate that parametric resonance sourced by AD condensates in such models naturally produces SGWB within reach of upcoming GW experiments.
+This study uncovers a generic class of well-motivated sources of SGWB and adds a timely complement to the rapidly growing field of probing particle physics with GWs, building on the ground-breaking discoveries of GWs by LIGO-Virgo-KAGRA collaboration Abbott et al.
 (2016a, b); Acernese et al.
-(2015) ushers in a new era of GW astronomy, followed by the more recent compelling evidence of SGWB identified by pulsar timing array (PTA) observations Carilli and Rawlings (2004); Weltman et al.
+(2015) and the recent strong evidence for SGWBs revealed by pulsar timing arrays Carilli and Rawlings (2004); Weltman et al.
 (2020); Lentati et al.
@@
 (2023a, b).
-This study uncovers a generic class of well-motivated sources of SGWB, making a timely complement to the literature.
-The study also reveals a new potential synergy between GW detection and laboratory experiments at energy, intensity, and neutrino frontiers, such as LHC, DUNE, SHiP, FASER Aad et al.
+Based on simple benchmark models, we show that successful baryogenesis requires new physics in the AD sector with a characteristic scale $m_{\Phi}={\cal O}(0.1-10)$ GeV, and the resulting GWs have a peak frequency in the range of ${\cal{O}}(1-100)$ Hz, within reach of upcoming detectors, including the Cosmic Explorer (CE) Abbott et al.
+(2017) and Einstein Telescope (ET) Hild et al.
+(2008), as well as DECIGO Kawamura et al.
+(2021) and Big Bang Observatory (BBO) Harry et al.
+(2006) (for the low-frequency tail).
+The study further uncovers a new potential complementarity between GW detection and laboratory experiments at energy, intensity, and neutrino frontiers, such as the LHC, DUNE, SHiP, and FASER Aad et al.
 (2021); Aaboud et al.
@@
 (2018); Barouki et al.
-(2022): interactions between the AD scalar $\Phi$ and the SM particles are necessary to transfer the asymmetry from the $\Phi$ field to SM baryons, while the characteristic new physics scale in the AD scalar sector is of ${\cal O}(0.1)$${\cal O}(10)$ GeV for viable models with detectable GW frequencies.
-The rest of the Letter is organized as follows: in the next section, we present the benchmark models and the prediction for the baryon asymmetry; following that, we show the results of lattice simulation for GW signals from these models.
-Finally, we conclude with an outlook.
-Benchmark Models.
-I. AD scalars: background dynamics and asymmetry generation
-We consider a complex scalar field $\Phi$, which is a SM gauge singlet with canonical kinetic terms and Einstein gravity. We work with a simple renormalizable polynomial potential:
-$V(\Phi)=\lambda_{\Phi}\left|\Phi\right|^{4}+m_{\Phi}^{2}\left|\Phi\right|^{2}-A(\Phi^{n}+\Phi_{\ast}^{n}),$ (1)
-with $n$ in principle being any integer $\leq 4$.
-We choose $n=2$ for concreteness; however, the results are not expected to change significantly with other choices for $n$.
-The first two terms manifest a $U(1)$ symmetry, which is broken by the third term as necessary to generate the $\Phi$-$\Phi^{\dagger}$ asymmetry, which later transfers to the baryon asymmetry.
-We consider the general possibility that $\Phi$, as a light spectator field during inflation, starts with a very large (even Planckian) field displacement Starobinsky and Yokoyama (1994); Cui et al.
+(2022), in search for new physics, depending on the details of the asymmetry transfer interactions between the AD field and the SM.
+Benchmark Models. I. AD scalars: background dynamics and asymmetry generation
+We consider a complex scalar field $\Phi$ (SM gauge singlet) with canonical kinetic terms, Einstein gravity, and a simple renormalizable polynomial potential:
+$V(\Phi)=\lambda_{\Phi}\left|\Phi\right|^{4}+m_{\Phi}^{2}\left|\Phi\right|^{2}-A(\Phi^{n}+\Phi^{\ast n}),$ (1)
+where $n\leq 4$ is an integer.
+We choose $n=2$ for concreteness.
+Although other choices may lead to similar results, $n=2$ stands out as a simple and natural option   [^2] For example, with the alternative choice of $n=4$, the generated asymmetry quickly freezes in at a large value, requiring a very small value of A to reproduce the observable baryon asymmetryRubakov and Gorbunov (2017); Hertzberg and Karouby (2014).
+In contrast, with a quadratic symmetry-breaking term (n = 2), the oscillations of the $\Phi$-condensate around the origin continue to generate and cancel asymmetry over time, partially suppressing the initially large asymmetry .
+The first two terms manifest a $U(1)$ symmetry, which is broken by the third term as necessary to generate the $\Phi$-$\Phi^{\dagger}$ asymmetry, precursor to the baryon asymmetry.
+We consider the general possibility that $\Phi$, as a light spectator during inflation, starts with a large (even Planckian) field displacement Starobinsky and Yokoyama (1994); Cui et al.
 (2024); Hardwick et al.
 (2017); Markkanen et al.
-(2018, 2019), e.g., simply by initial condition, or by quantum fluctuation in the inflationary de Sitter space.
-We have taken into account the isocurvature constraints from CMB data, which is relevant to light spectator fields in general, and found that they can be generally satisfied (see Section .2 of the Supplemental Material for details).
-We denote the typical value at the end of inflation as $\Phi_{\mathrm{in}}$, the initial value for our calculations  [^2] Other mechanisms can be invoked to drive the $\Phi$ field to large values during inflation, like a non-minimal coupling to gravity or a direct coupling to the inflaton.
-Our results for the GWs and baryon asymmetry do not depend on the origin of $\Phi_{\mathrm{in}}$..
-As long as the (radial) field is light, $H\gg H_{\mathrm{osc}}\simeq\sqrt{V_{,|\Phi|}/|\Phi|}$, where $H_{\mathrm{osc}}$ is the effective mass of $\Phi$ Cui et al.
+(2018, 2019), for instance, due to de Sitter fluctuations during inflation.
+We have verified that the isocurvature constraints from CMB data, which are relevant for any light spectator field, can be satisfied.We denote the typical value at the end of inflation as $\Phi_{\mathrm{in}}$, which is the initial value for our calculations   [^3] Other mechanisms can drive the $\Phi$ field to large values during inflation, like a non-minimal coupling to gravity or a direct coupling to the inflaton.
+Our results for GWs and $n_{b}-n_{\bar{b}}$ do not depend on the origin of $\Phi_{\mathrm{in}}$.
+.
+As long as the (radial) field is light, $H\gg H_{\mathrm{osc}}\simeq\sqrt{V_{,|\Phi|}/|\Phi|}$, where $H_{\mathrm{osc}}$ is the effective mass of $|\Phi|$ Cui et al.
 (2024), it remains frozen  [^4] Since we consider small $U(1)$ symmetry-breaking terms, the angular field will always be light when the radial field is light..
-The initial phase angle of $\Phi$, $\theta=Arg(\Phi)$, is also set as $\theta_{\mathrm{in}}$ at the end of inflation in our observable patch of the universe.
+The initial phase of $\Phi$, $\theta=Arg(\Phi)$, is also set as $\theta_{\mathrm{in}}$ at the end of inflation in our observable patch of the Universe.
 As the Universe expands, the Hubble rate drops, and at $H\simeq H_{\mathrm{osc}}$ the field starts rolling.
 For our parameters, the rolling starts during the radiation-dominated (RD) era, well after the end of inflation.
-We analyze the motion of $\Phi$ in terms of its real and imaginary parts (as defined in $\Phi={1\over\sqrt{2}}(\phi_{R}+i\phi_{I})$), which obey the following equations of motion:
+The real and imaginary parts of $\Phi\equiv\frac{1}{\sqrt{2}}(\phi_{R}+i\phi_{I})$ obey the following equations of motion:
 $\ddot{\phi}_{{}_{\mathrm{R}}}$ $+3H\dot{\phi}_{{}_{\mathrm{R}}}+m_{{}_{\mathrm{R}}}^{2}\phi_{{}_{\mathrm{R}}}+\lambda_{\Phi}(\phi_{{}_{\mathrm{R}}}^{2}+\phi_{{}_{\mathrm{I}}}^{2})\phi_{{}_{\mathrm{R}}}=0,$ (2)
@@
 where $m^{2}_{R,I}=m_{\Phi}^{2}\mp 2A$.
-The Hubble expansion rate in a RD Universe is $H\propto t^{1/2}$.
-The overdots denote derivatives with respect to cosmic time $t$.
-We see that the two components have different masses in the presence of $A\neq 0$, and thus, the trajectory in the $\phi_{R}$-$\phi_{I}$ plane is expected to resemble a shrinking ellipse instead of a straight line, even for vanishing initial velocities.
-The background field evolution, neglecting the symmetry-breaking term ($A\ll m_{\Phi}^{2}$), is initially dominated by the $\left|\Phi\right|^{4}$ term when the amplitude scales as $\phi\propto 1/a$.
-When the amplitude drops below $|\Phi_{\ast}|=m_{\Phi}/\sqrt{\lambda_{\Phi}}$ at $t=t_{*}$, the $\left|\Phi\right|^{2}$ term dominates the oscillation and the fields evolve approximately as $\phi_{R,I}\sim a^{-3/2}\cos(m_{R,I}t+c_{R,I})$ (see Section .3 of the Supplemental Material for details).
-The $\Phi$-$\Phi^{\dagger}$ asymmetry, which will be transferred later to the baryon (or lepton) number, is $n_{\Phi}(t)=i\left(\Phi^{\dagger}\dot{\Phi}-\dot{\Phi}^{\dagger}\Phi\right)=\dot{\phi}_{\mathrm{R}}\phi_{{}_{\mathrm{I}}}-\phi_{{}_{\mathrm{R}}}\dot{\phi}_{\mathrm{I}}$.
-$n_{\Phi}$ directly corresponds to $n_{B}$ or $n_{L}$ in the absence of significant washout (see discussion later).
-Using the above approximations for $\phi_{R,I}$ and including the terms $\Gamma_{\Phi}\dot{\phi}_{R,I}$ to model the decay of $\Phi$ to SM states, we can write the co-moving baryon asymmetry generated during the time $t>t_{*}$ as Rubakov and Gorbunov (2017); Lloyd-Stubbs and McDonald (2021, 2023); Mohapatra and Okada (2021):
+The Hubble expansion rate in an RD Universe is $H\propto t^{1/2}$.
+An overdot denotes derivative with respect to cosmic time $t$.
+If $A\neq 0$, the real and imaginary components have different masses, so their trajectories in the $\phi_{R}$-$\phi_{I}$ plane resemble shrinking ellipses (not straight lines), even for zero initial velocities.
+The background field evolution (assuming $A\ll m_{\Phi}^{2}$) is initially dominated by the $\left|\Phi\right|^{4}$ term, where the amplitude scales as $\phi\propto 1/a$.
+When the amplitude drops below $|\Phi_{\ast}|=m_{\Phi}/\sqrt{\lambda_{\Phi}}$ at $t=t_{*}$, the $\left|\Phi\right|^{2}$ term dominates the oscillation and the fields evolve approximately as $\phi_{R,I}\sim a^{-3/2}\cos(m_{R,I}t+c_{R,I})$, with $c_{R,I}$ being a constant phase (see Supplemental Material, Section C, for details).
+The $\Phi$-$\Phi^{\dagger}$ asymmetry, which will be subsequently transferred to the baryon (or lepton) number, is $n_{\Phi}(t)\equiv i\left(\Phi^{\dagger}\dot{\Phi}-\dot{\Phi}^{\dagger}\Phi\right)=\dot{\phi}_{\mathrm{R}}\phi_{{}_{\mathrm{I}}}-\phi_{{}_{\mathrm{R}}}\dot{\phi}_{\mathrm{I}}$.
+In the absence of significant washout, $n_{\Phi}$ directly corresponds to $n_{B}$ or $n_{L}$ (see the Discussion later).
+Using the above approximations for $\phi_{R,I}$ and including the terms $\Gamma_{\Phi}\dot{\phi}_{R,I}$ to model the decay of $\Phi$ to SM states (the specific form of $\Gamma_{\Phi}$ depends on the decay channels, which we elaborate on later), the comoving baryon asymmetry generated during the time $t>t_{*}$ is Rubakov and Gorbunov (2017); Lloyd-Stubbs and McDonald (2021, 2023); Mohapatra and Okada (2021):
 $\left(\frac{a(t)}{a_{\mathrm{in}}}\right)^{3}n_{B}(t)\simeq 4A\phi_{{}_{\mathrm{R,in}}}\phi_{{}_{\mathrm{I,in}}}\left(\frac{\phi_{\mathrm{in}}}{\phi_{*}}\right)\times$
-[…]
-where the exponential factor accounts for the decrease in $\Phi$ amplitude due to its decay to SM particles with a constant decay rate $\Gamma_{\Phi}$, enabling the asymmetry transfer to the SM baryons.
-The specific form of $\Gamma_{\Phi}$ depends on the decay channels, which we will elaborate on later.
+$\int_{t_{*}}^{t}dt^{\prime}\cos{(m_{{}_{\mathrm{R}}}(t^{\prime}-t_{\ast}))}\sin{(m_{{}_{\mathrm{I}}}(t^{\prime}-t_{\ast}))}e^{-\Gamma_{\Phi}(t^{\prime}-t_{*})},$ (4)
+which can be further simplified for $A\ll m_{\phi}^{2}$ (the detailed derivation is presented as Supplemental Material) and scales as
+$\frac{a^{3}(t)}{a^{3}_{\mathrm{in}}}n_{B}(t)\propto\int_{t_{*}}^{t}dt^{\prime}\sin\left(\frac{2A}{m_{\Phi}}(t-t_{*})\right)e^{-\Gamma_{\Phi}(t-t_{*})}\,.$ (5)
 Evaluating the above integral for $t\rightarrow\infty$ and factoring in the expansion during the RD Universe, we express $n_{B}/s$ as
-$\frac{n_{B}}{s}=\begin{cases}\left(\frac{4\alpha^{3}}{\lambda_{\Phi}k_{T_{d}}^{6}}\right)^{1/4}\epsilon_{\Phi}\frac{m_{\Phi}M_{\mathrm{Pl}}}{T_{d}^{2}}\sin(2\theta);&\gamma_{\Phi}\gg 2\epsilon_{\Phi}\\ \left(\frac{\alpha^{3}k_{T_{d}}^{2}}{64\lambda_{\Phi}}\right)^{1/4}\frac{1}{\epsilon_{\Phi}}\frac{T_{d}^{2}}{m_{\Phi}M_{\mathrm{P}}}\sin(2\theta);&\gamma_{\Phi}\ll 2\epsilon_{\Phi}\end{cases}$ (5)
+$\frac{n_{B}}{s}=\begin{cases}\left(\frac{4\alpha^{3}}{\lambda_{\Phi}k_{T_{d}}^{6}}\right)^{1/4}\epsilon_{\Phi}\frac{m_{\Phi}M_{\mathrm{Pl}}}{T_{d}^{2}}\sin(2\theta);&\gamma_{\Phi}\gg 2\epsilon_{\Phi}\\ \left(\frac{\alpha^{3}k_{T_{d}}^{2}}{64\lambda_{\Phi}}\right)^{1/4}\frac{1}{\epsilon_{\Phi}}\frac{T_{d}^{2}}{m_{\Phi}M_{\mathrm{Pl}}}\sin(2\theta);&\gamma_{\Phi}\ll 2\epsilon_{\Phi}\end{cases}$ (6)
 where $M_{\mathrm{Pl}}\simeq 2.4\times 10^{18}$ GeV is the reduced Planck mass, and we define the dimensionless quantities: $\epsilon_{\Phi}=A/m_{\Phi}^{2}$, $\gamma_{\phi}=\Gamma_{\Phi}/m_{\Phi}$ and $k_{T_{d}}=\sqrt{\pi^{2}g(T_{d})/90}$.
-$T_{d}$ is the temperature of the Universe when the $\Phi$ decay is completed at $H(T_{d})=\Gamma_{\Phi}$, $\alpha$ is the initial fraction of energy density in the spectator sector $\Phi$ relative to the total energy density, and $\theta=\theta_{\mathrm{in}}$ is defined earlier.
-In case the asymmetry is first transferred to SM leptons, a redistribution factor of $28/79$ due to the EW sphaleron would be added in the RHS of Eq.
+$T_{d}$ is the temperature of the Universe characterizing $\Phi$ decay, as defined by $H(T_{d})=\Gamma_{\Phi}$, $\alpha$ is the initial fraction of energy density in the spectator sector $\Phi$ relative to the total energy density, and $\theta=\theta_{\mathrm{in}}$ is defined earlier.
+The exact form of $n_{B}/s$, from which Eq.
+(6) is derived, is given as Supplemental Material.
+The change in the scaling of the final asymmetry with the normalized symmetry–breaking parameter $\epsilon_{\Phi}\equiv A/m_{\Phi}^{2}$ shown by the two branches of Eq.
+(6) can be understood simply in terms of the competition between two timescales: the oscillation period of the asymmetry and the lifetime of the scalar.
+The symmetry–breaking parameter $\epsilon_{\Phi}$ induces a mass splitting between the particle and antiparticle components of $\Phi$, leading to an oscillation of the charge asymmetry with frequency $\omega_{\Phi}\sim\Delta m_{\Phi}\sim\epsilon_{\Phi}m_{\Phi}$ and period $T_{\mathrm{asy}}\sim 1/(\epsilon_{\Phi}m_{\Phi})$.
+At the same time, the scalar decays with rate $\Gamma_{\Phi}$, corresponding to a lifetime $\tau_{\Phi}\sim 1/\Gamma_{\Phi}$.
+The net asymmetry transferred to the thermal bath is proportional to the time integral of an oscillatory source with an exponential envelope, schematically $n_{B}\propto\int_{0}^{\infty}dte^{-\Gamma_{\Phi}t}\sin(\omega_{\Phi}t)$, so the final result depends on whether the decay occurs before or after many oscillations.
+In the regime $\Gamma_{\Phi}\gg\omega_{\Phi}$ (i.e.
+$\tau_{\Phi}\ll T_{\mathrm{asy}}$), the integral is dominated by very early times, when the sine function has not yet completed a full oscillation and may be approximated as $\sin(\omega_{\Phi}t)\simeq\omega_{\Phi}t$.
+The generated asymmetry is then essentially fixed before sign–flipping oscillations set in, and scales linearly with the oscillation frequency (equivalently with the asymmetry parameter $\epsilon_{\Phi}$) as
+$n_{B}\propto\int_{0}^{\sim 1/\Gamma_{\Phi}}dte^{-\Gamma_{\Phi}t}\omega_{\Phi}t\sim\frac{\omega_{\Phi}}{\Gamma_{\Phi}}\propto\epsilon_{\Phi}.$ (7)
+By contrast, in the regime $\Gamma_{\Phi}\ll\omega_{\Phi}$ (i.e.
+$\tau_{\Phi}\gg T_{\mathrm{asy}}$), the source oscillates many times before the field decays, spending comparable time with positive and negative signs.
+As a result, contributions from successive oscillations partially cancel, and the net asymmetry is a residue arising from the slow decay of the envelope.
+To see the overall scale of the asymmetry for this limit, we note that the contribution from a single half-period of the oscillation is of order $\int dt\sin(\epsilon_{\Phi}m_{\Phi}t)\sim 1/(\epsilon_{\Phi}m_{\Phi})$, which sets the characteristic scale of the integral in Eq.
 (5).
-Eq.
-(5) is derived assuming that the asymmetry is generated when the potential is dominated by the $|\Phi|^{2}$ term; thus, we require that the background dynamics prior to $t_{*}$, defined as when $|\Phi(t_{\ast})|=m_{\Phi}/\sqrt{\lambda_{\Phi}}$, is not affected by the $A$ term.
-The corresponding condition on $A$ can be estimated as follows: The only term in the potential that involves the angular degree of freedom is the symmetry-breaking term $2A|\Phi(t)|^{2}\cos 2\theta(t)\subset V(|\Phi|,\theta)$, indicating an upper bound on the effective mass along the angular direction: $|m^{2}_{\theta}|\leq 2A$.
-The angular degree of freedom is frozen at $\theta_{\mathrm{in}}$ for $t<t_{*}$, if $|m^{2}_{\theta}|<H^{2}(t_{*})$ Lloyd-Stubbs and McDonald (2023), which translates to $\epsilon_{\Phi}\equiv A/m_{\Phi}^{2}<({\mu^{2}}/{16\alpha})\left({\phi_{\mathrm{in}}}^{2}/{M_{\mathrm{Pl}}}^{2}\right)$, where $\mu^{2}\equiv m_{\Phi}^{2}/(\lambda_{\Phi}\phi_{\mathrm{in}}^{2})$, and $\alpha$ is defined below Eq.
-5.
-As we see in the next section, typically ${\phi_{\mathrm{in}}}\simeq M_{\mathrm{Pl}}$ and $\mu^{2}\lesssim 10^{-5}$ are required for detectable GW signals.
-With larger $A$, motion along the angular direction is non-negligible, and $\theta(t)$ would have already evolved towards its minimum before $t=t_{*}$, thus reducing $\sin\theta$ and suppressing the asymmetry (see Eq.
-(5)).
+Including all subsequent oscillations leads to an alternating geometric series, which can be summed and consequently multiplies the result of the first oscillation by $1/2$ for the case with $\gamma_{\phi}/2\epsilon_{\Phi}\ll 1$.
+The final asymmetry thus scales as
+$n_{B}\propto$ $\int_{0}^{\infty}dt\sin(\epsilon_{\Phi}m_{\Phi}t)$
+$\simeq$ $\frac{1}{2}\times\int_{0}^{\pi/(2\epsilon_{\Phi}m_{\Phi}))}dt\sin(\epsilon_{\Phi}m_{\Phi}t)\sim\frac{1}{2}\frac{1}{\epsilon_{\Phi}m_{\Phi}}.$ (8)
+In summary, a quadratic symmetry-breaking term induces a constant mass splitting between the real and imaginary components of $\Phi$, leading to a single well–defined oscillation frequency $\omega_{\Phi}\propto\epsilon_{\Phi}m_{\Phi}$.
+This simple harmonic structure makes the final asymmetry an explicit function of $\omega_{\Phi}$, and hence of $\epsilon_{\Phi}$, producing the characteristic two-regime behavior [^5] Note that the $\epsilon_{\Phi}\rightarrow 0$ limit is only reached through the upper branch of Eq.
+(6), leading to $n_{B}/s\rightarrow 0$ as expected.
+$n_{B}\propto\epsilon_{\Phi}$ for $\Gamma_{\Phi}\gg\omega_{\Phi}$ and $n_{B}\propto 1/\epsilon_{\Phi}$ for $\Gamma_{\Phi}\ll\omega_{\Phi}$, shown in Eq.
+(6).
+In the case where the asymmetry in the $\Phi$ field is first transferred to an SM lepton asymmetry $n_{L}$, which is later converted to a baryon asymmetry $n_{B}$ via the sphaleron process, the above result gets multiplied by a factor of $28/79$ since Harvey and Turner (1990); Asaka et al.
+(2002); Garcia-Bellido et al.
+(1999) $n_{B}/s=-(28/79)\times n_{L}/s\,.$
+Equation (6) is derived assuming that the asymmetry is generated when the potential is dominated by the $|\Phi|^{2}$ term, thus requiring that the background dynamics prior to $t_{*}$ is not affected by the $A$ term.
+The corresponding condition on $A$ is derived as follows: the only term in the potential that involves the angular degree of freedom is the symmetry-breaking term $2A|\Phi(t)|^{2}\cos 2\theta(t)\subset V(|\Phi|,\theta)$, indicating an upper bound on the effective mass along the angular direction: $|m^{2}_{\theta}|\leq 2A$.
+The angular degree of freedom is frozen at $\theta_{\mathrm{in}}$ for $t<t_{*}$ if $|m^{2}_{\theta}|<H^{2}(t_{*})$ Lloyd-Stubbs and McDonald (2023), which translates into $\epsilon_{\Phi}\equiv A/m_{\Phi}^{2}<({\mu^{2}}/{16\alpha})\left(\phi_{\mathrm{in}}^{2}/M_{\mathrm{Pl}}^{2}\right)$, where $\mu^{2}\equiv m_{\Phi}^{2}/(\lambda_{\Phi}\phi_{\mathrm{in}}^{2})$, and $\alpha$ is defined below Eq.
+6.
+As we see in the next section, ${\phi_{\mathrm{in}}}\simeq M_{\mathrm{Pl}}$ and $\mu^{2}\lesssim 10^{-5}$ are required for detectable GW signals.
+With larger $A$, motion along the angular direction is non-negligible, and $\theta(t)$ would have already evolved toward its minimum before $t=t_{*}$, thus reducing $\sin\theta$ and suppressing the asymmetry [see Eq.
+(6)].
 For our benchmark examples, the chosen $A$ values satisfy this self-consistency condition.
@@
 For $\epsilon_{\Phi}=10^{-20}$ the observed $n_{B}/s\sim 10^{-10}$ can be realized with $T_{d}\sim 2.5\times 10^{7}$ GeV when $m_{\Phi}=10\,\mathrm{GeV}$.
-As we will discuss, transfer of the asymmetry to the baryon sector may involve significant washout (model-dependent), requiring a larger initial $\Phi$ asymmetry to compensate.
-Therefore we also consider the parameter points that lead to $n_{\Phi}/s\gg 10^{-10}$, for example, with the same $\epsilon_{\Phi}$, but $T_{d}\sim 1$ GeV, one obtains $n_{\Phi}/s\sim 4\times 10^{-5}$.
-A more detailed parameter scan is given in Section .3 of the Supplemental Material and Fig.
-S1, demonstrating the possibility of producing a large initial asymmetry to compensate for potential (model-dependent) washout effects.
-II. Transfer to SM baryon asymmetry.
+As we will discuss, transferring the asymmetry to the baryon sector may involve a significant (model-dependent) washout, requiring a larger initial $\Phi$ asymmetry to compensate.
+Therefore, we also consider parameter points that lead to $n_{\Phi}/s\gg 10^{-10}$, for example, the same $\epsilon_{\Phi}$ with $T_{d}\sim 1$ GeV gives $n_{\Phi}/s\sim 4\times 10^{-5}$.
+A more detailed parameter scan is provided in Section C of Supplemental Material, demonstrating the potential to produce a large initial asymmetry that compensates for possible washout effects.
+II. Transfer to the SM baryon asymmetry .
 The asymmetry in the $\Phi$ field generated during its oscillation needs to eventually be transferred to the SM baryon asymmetry via its subsequent decay to baryons or leptons (further assisted by electroweak sphalerons in the latter case).
-This can generally be realized by a B- or L-violating coupling involving $\Phi$ in terms of effective higher dimensional operators, such as $\Phi QQQL/\Lambda^{3}$ Allahverdi and Mazumdar (2012), $\Phi UDDUDD/\Lambda^{6}$ and $\Phi LLELLE/\Lambda^{6}$.
+This can generally be realized by a B- or L-violating coupling involving $\Phi$ in terms of effective higher-dimensional operators, such as $\Phi QQQL/\Lambda^{3}$ Allahverdi and Mazumdar (2012), $\Phi UDDUDD/\Lambda^{6}$, and $\Phi LLELLE/\Lambda^{6}$.
 Whether the related constraints, for instance, from $n-\bar{n}$ oscillation, can be satisfied depends on the specifics of the UV completion, including flavor structure in the couplings.
-As a proof of principle, in the following, we outline examples of renormalizable UV complete models realizing the $\Phi$ to baryon asymmetry transfer and summarize the conditions on viable model parameters.
-Further details of the analysis are given in Section .4 of the Supplemental Material.
+As a proof of principle, we outline examples of renormalizable, UV-complete models that realize the $\Phi$ to baryon asymmetry transfer in the following, and summarize the conditions on viable model parameters.
+Further details are given in Section D of Supplemental Material.
 The core results of this work are independent of the details of the asymmetry transfer mechanism.
@@
 $\mathcal{L}\supset y_{N,~i}\Phi\bar{N_{i}}^{C}{N_{i}}+g_{\nu,~ij}\bar{N}_{i}HL_{j}+{\mathrm{h.c.}},$ (9)
-where $N_{i}$’s are right-handed (RH) Majorana neutrinos (with $i=1,2,3$ as a typical choice), $H$ is SM Higgs, $L_{j}$ are SM lepton doublets.
-This model potentially relates to neutrino mass generation, which nevertheless is not imposed as a requirement here.
+where $N_{i}$’s are right-handed Majorana neutrinos (with $i=1,2,3$ as a typical choice), $H$ is SM Higgs, and $L_{j}$ are SM lepton doublets.
+This model potentially relates to neutrino mass generation, which, nevertheless, is not imposed as a requirement here.
 We consider the simple scenario where $\Phi$ decays to $NN$, then $N$ transfers the asymmetry to SM leptons via freeze-out or freeze-in processes enabled by the $NHL$ coupling, as considered in the literature of leptogenesis Akhmedov et al.
@@
 (B) $\Phi$ decay to baryons (direct baryogenesis)
-A simple renormalizable model, in this case, includes:
+A simple renormalizable model, in this case, includes
 $\mathcal{L}\supset y_{\chi}\Phi\chi\chi+g_{i}\chi\bar{u}_{i}\varphi+\lambda_{ij}\varphi d_{i}d_{j}+{\mathrm{H.c.}},$ (10)
-where $\chi$ is a singlet Majorana fermion, $\varphi$ is an up-type diquark scalar, and the flavor indices $i,~{}j=1,2,3$ in down-type quark combination $d_{i}d_{j}$ should be antisymmetric.
+where $\chi$ is a singlet Majorana fermion, $\varphi$ is an up-type diquark scalar, and the flavor indices i,j=1,2,3i,~j=1,2,3 in down-type quark combination $d_{i}d_{j}$ should be antisymmetric.
 In the viable parameter region, the decay proceeds with $\Phi\rightarrow\chi\chi$, followed by $\chi\rightarrow udd$.
-Both models are subject to cosmological conditions which impose lower limits on $m_{\Phi}$ and $m_{\chi}$ or $m_{N}$.
+Both models are subject to cosmological conditions that impose lower limits on $m_{\Phi}$, $m_{\chi}$, or $m_{N}$.
 For $\Phi$ to decay to baryons (before BBN), it is subject to kinematic conditions: $m_{\Phi}\gtrsim 2$ GeV, and $m_{\chi}\gtrsim 1$ GeV.
-For $\Phi$ asymmetry to be transferred to the SM leptons via $N$ before the EW phase transition, $m_{\Phi}>2m_{N}\gtrsim O(0.1)$ GeV Klarić et al.
+For $\Phi$ asymmetry to be transferred to SM leptons via $N$ before the EW phase transition requires $m_{\Phi}>2m_{N}\gtrsim O(0.1)$ GeV Klarić et al.
 (2021a).
-In addition, for both models, the potential asymmetry washout effects should be either suppressed or can be compensated by a larger initial $\Phi$ asymmetry produced during its oscillation (which has been shown viable in the last section).
-Furthermore, relevant laboratory experiments at energy and intensity frontiers searching for B-violation, color-charged new particles, sterile neutrinos, etc., may provide additional constraints on model parameters.
-In Section .4 of the Supplemental Material, we elaborate on relevant constraints and demonstrate examples of viable parameter regions.
-On the other hand, as we will see in our GW results, an upper limit in mass of $m_{\Phi}\lesssim O(10)$ GeV corresponds to the currently detectable GW peak frequency range, which is below $\sim O(100)$ Hz.
-This, combined with the aforementioned cosmological lower limit of $m_{\Phi}\gtrsim O(0.1-1)$ GeV, demonstrates that intriguingly the parameter range of our interest points to new physics at the energy scale of $O(0.1-10)$ GeV, indicating a new potential complementarity between GW detection and laboratory searches across particle physics frontiers.
+In addition, for both models, the potential asymmetry washout effects should be either suppressed or compensated by a larger initial $\Phi$ asymmetry produced during its oscillation (which has been shown to be viable in the last section).
+Furthermore, relevant laboratory experiments at the energy and intensity frontiers, searching for B-violation, color-charged new particles, and sterile neutrinos, among others, may provide additional constraints on model parameters.
+In Section D of Supplemental Material, we elaborate on relevant constraints and demonstrate examples of viable parameter regions.
 Parametric Resonance and GW Signals.
 We have seen that the generation of the baryon asymmetry is determined by the classical motion of $\Phi$ as an averaged background field.
-In addition to this classical motion of $\Phi$, we must consider the corresponding fluctuations around it, which leads to parametric resonance that may source GWs.
-We proceed by decomposing $\phi_{R,I}=\bar{\phi}_{R,I}(t)+\delta\phi_{R,I}(x,t)$, where $\bar{\phi}_{R,I}$ describe the classical background motion and $\delta\phi_{R,I}(x,t)$ correspond to the fluctuations, which have a quantum origin.
+In addition to this classical motion of $\Phi$, the corresponding fluctuations around it can lead to parametric resonance and source GWs.
+We decompose $\phi_{R,I}=\bar{\phi}_{R,I}(t)+\delta\phi_{R,I}(x,t)$, where $\bar{\phi}_{R,I}$ describe the classical background motion and $\delta\phi_{R,I}(x,t)$ correspond to the fluctuations, which have a quantum origin.
 The evolution of background fields $\bar{\phi}_{R,I}$ follows Eqs.
-(2), (3), whereas we expand the fluctuations in Fourier modes as
+(2) and (3), whereas we expand the fluctuations in Fourier modes as
 $\delta\phi_{R}(x,t)=\int{d^{3}k\over(2\pi)^{3}}\left[a_{R,k}e^{i\vec{k}\cdot\vec{x}}\delta\phi_{R,k}(t)+H.c.\right]$ (11)
 and similarly for $\delta\phi_{I}(x,t)$.
-As the background fields oscillate around the minimum of the potential, the fluctuations in momentum-space evolve according to
+As the background fields oscillate around the minimum of the potential, the fluctuations in momentum space evolve according to
 $\ddot{\delta\phi}_{R,I}+3H\dot{\delta\phi}_{R,I}+m_{R,I}^{2}{\delta\phi_{R,I}}+{k^{2}\over a^{2}}{\delta\phi_{R,I}}$
-$+\lambda_{\phi}(3\bar{\phi}_{R,I}^{2}+\bar{\phi}_{I,R}^{2})\delta\phi_{R,I}+2\bar{\phi}_{I}\bar{\phi}_{R}\delta\phi_{I,R}=0$ (9)
+$\quad+\lambda_{\phi}(3\bar{\phi}_{R,I}^{2}+\bar{\phi}_{I,R}^{2})\delta\phi_{R,I}+2\bar{\phi}_{I}\bar{\phi}_{R}\delta\phi_{I,R}=0.$ (12)
 We thus see that, in general, the equations for $\delta\phi_{R,I}$ are coupled.
 We first consider the representative limit of $A=0$, where $m_{R}=m_{I}=m_{\Phi}$ and the background field trajectory is simply a straight line through the origin (for vanishing initial velocities, as is the case here).
-In this case, we can rotate our field basis and define fields in the radial direction (along the background motion) ϕsubscriptitalic-$\phi_{\parallel}$ and perpendicular to it (orthogonal direction) as ϕsubscriptitalic-$\phi_{\perp}$.
-By construction, $\bar{\phi}_{\perp}=0$, and the fluctuation equations for ϕsubscriptitalic-$\phi_{\parallel}$ and ϕsubscriptitalic-$\phi_{\perp}$ decouple, namely,
-$\ddot{\delta\phi}_{\parallel,\perp}+3H\dot{\delta\phi}_{\parallel,\perp}+\left({k^{2}\over a^{2}}+m_{\Phi}^{2}+q_{\parallel,\perp}\lambda_{\Phi}\bar{\phi}_{\parallel,\perp}^{2}\right)\delta\phi_{\parallel,\perp}=0$ (10)
+In this case, we can rotate our field basis and define fields in the radial direction (along the background motion) $\phi_{\parallel}$ and perpendicular to it (orthogonal direction) as $\phi_{\perp}$.
+By construction, $\bar{\phi}_{\perp}=0$, and the fluctuation equations for $\phi_{\parallel}$ and $\phi_{\perp}$ decouple, namely,
+$\ddot{\delta\phi}_{\parallel,\perp}+3H\dot{\delta\phi}_{\parallel,\perp}+{k^{2}\over a^{2}}{\delta\phi}_{\parallel,\perp}+m_{\Phi}^{2}{\delta\phi}_{\parallel,\perp}$
+$+q_{\parallel,\perp}\lambda_{\Phi}\bar{\phi}_{\parallel,\perp}^{2}\delta\phi_{\parallel,\perp}=0,$ (13)
 where $q_{\parallel,\perp}=\{3,1\}$.
-As the background field oscillates quasi-periodically around the origin, fluctuations can be exponentially amplified for certain ranges of wavenumbers.
-Through the exponential amplification, the occupation number of these modes grows rapidly, and they can thus be treated classically Polarski and Starobinsky (1992); Kofman et al.
+Quasiperiodic background-field oscillations can exponentially amplify fluctuations within certain wave-number ranges.
+This rapid growth of mode occupation numbers allows these modes to be treated classically Polarski and Starobinsky (1992); Kofman et al.
 (1997).
-This phenomenon is referred to as “preheating” Greene et al.
-(1997) in the context of primordial particle production following inflation, while it also can occur for oscillating spectator scalar fields in general (see, for instance, our related work Cui et al.
+This phenomenon, originally studied for exponential particle production (“preheating”) after inflation Greene et al.
+(1997), can also occur for any oscillating scalar (see, for instance, Cui et al.
 (2024)).
-To quantify parametric resonance, we start by defining a dimensionless field amplitude $\tilde{\phi}=\phi/\phi_{\mathrm{in}}$ and a corresponding conformal scaling for time as $\mathrm{d}\eta=\sqrt{\lambda_{\Phi}}\phi_{\mathrm{in}}\mathrm{d}t/a$ with $\phi_{\mathrm{in}}$ being the initial amplitude of the scalar field $\phi_{\mathrm{in}}=|\Phi|_{\mathrm{in}}$.
-The dimensionless fluctuation equation becomes (derivatives are over $\eta$)
-${\delta\phi}_{\parallel,\perp}^{\prime\prime}+2\tilde{H}{\delta\phi}_{\parallel,\perp}^{\prime}+\left({\tilde{k}^{2}}+a^{2}\mu^{2}+q_{\parallel,\perp}a^{2}\tilde{\phi}^{2}\right)\delta\phi_{\parallel,\perp}=0$ (11)
+To quantify parametric resonance, we define a dimensionless field amplitude $\tilde{\phi}=\phi/\phi_{\mathrm{in}}$ and a corresponding conformal scaling for time as $\mathrm{d}\eta=\sqrt{\lambda_{\Phi}}\phi_{\mathrm{in}}\mathrm{d}t/a$ with $\phi_{\mathrm{in}}$ being the initial amplitude of the scalar field $\phi_{\mathrm{in}}=\left|\Phi_{\mathrm{in}}\right|$.
+The dimensionless fluctuation equation becomes ($\delta\phi^{\prime}\equiv d\delta\phi/d\eta$):
+${\delta\phi}_{\parallel,\perp}^{\prime\prime}+2\tilde{H}{\delta\phi}_{\parallel,\perp}^{\prime}+\left({\tilde{k}^{2}}+a^{2}\mu^{2}+q_{\parallel,\perp}a^{2}\tilde{\phi}^{2}\right)\delta\phi_{\parallel,\perp}=0,$ (14)
 where $\tilde{k}^{2}=k^{2}/\lambda\phi_{\mathrm{in}}^{2}$ and $\tilde{\phi}=\bar{\phi}/\phi_{\mathrm{in}}$.
-[…]
-The growth of fluctuations and their back-reaction break the homogeneous oscillation of $\Phi$, leading to the inhomogeneous structure of scalar fields with an effective anisotropic stress tensor $\Pi_{ij}^{TT}=\nabla_{i}\phi^{a}\nabla_{j}\phi^{a}$ (where the repeated index `$a$’ indicates the number of the scalar components and a summation is implied).
-These time-dependent field inhomogeneities source (sub-Hubble) gravitational waves whose frequency depends on the characteristic energy scale involved $H_{c}$.
-The peak amplitude and frequency of GW spectra generated from sources peaked at some characteristic scale $k_{c}$ in momentum-space can be estimated as Giblin and Thrane (2014)
-$f_{\mathrm{GW}}^{\mathrm{peak}}=$ $2.7\times 10^{10}\sqrt{\frac{H_{c}}{M_{\mathrm{P}}}}\frac{k_{c}}{H_{c}}\,{\mathrm{Hz}}\,,$ (12)
-$\Omega_{\mathrm{GW}}^{\mathrm{peak}}=$ $2.3\times 10^{-4}\,\alpha^{2}\,\beta\,w\left(\frac{k_{c}}{\sigma}\right)\left(\frac{H_{c}}{k_{c}}\right)^{2}.$ (13)
-where $\alpha$ is a measure of the fraction of the energy in the GW source here the scalar field inhomogeneities and this $\alpha$ coincides with the $\alpha$ we defined below Eq.
-5 relative to the total energy density of the Universe, and $\beta$ describes the degree of anisotropy of the source which must be determined from simulations.
-In typical scenarios $\beta=\mathcal{O}(0.01-0.1)$.
-The equation of the state of the Universe at the time of the production of GWs for a radiation-dominated Universe is $w=1/3$ and $\sigma$ is the width of the source in momentum-space, which can be taken as $\sigma\sim k_{c}$ for peaked sources (see Ref.
-Giblin and Thrane (2014) for the derivation and e.g.
-Refs.
-Cui and Sfakianakis (2023); Cui et al.
-(2024) for further applications of Eqs (12) and (13)).
+There are two essential resonance parameters: $\mu^{2}=\frac{m_{\Phi}^{2}}{\lambda_{\Phi}\phi_{\mathrm{in}}^{2}}$ and $q_{\parallel,\perp}$.
+As shown in Greene et al.
+(1997), a large $\mu^{2}$ can suppress the resonance and determine if the system displays stable resonance, stochastic resonance, or no decay.
+We choose to focus on $\mu\leq 10^{-5}$, such that the system is always in a stable resonance regime.
+For such small $\mu^{2}$, the system is similar to the well-studied case of a quartic field during preheating, whereas $q_{\parallel,\perp}$ is fixed in our case due to the rotational symmetry of the potential for $A=0$.
+Section E of Supplemental Material discusses in detail the effect of $q_{\parallel,\perp}$ on the growth of $\delta\phi_{\parallel,\perp}$ and shows that for sufficiently small values of the asymmetry parameter $A\lesssim 0.005m_{\Phi}^{2}$, the parametric resonance dynamics is indistinguishable from the case $A=0$, and thus the GW amplitude remains insensitive to the value of $A$.
+Fortunately, the parameter range of $A$ that produces sufficient baryon asymmetry, as considered in our benchmarks, is well within this regime.
+On a technical note, we use the parametrization of real and imaginary components of $\Phi$ ($\phi_{R}$, $\phi_{I}$) when computing the baryon asymmetry, whereas we are using the radial direction and its orthogonal ($\phi_{\parallel}$, $\phi_{\perp}$) when computing parametric resonance for $A\ll m_{\Phi}^{2}$.
+This is due to the fact that the baryon asymmetry scales with $A$ and thus the difference between the $\phi_{I}$ and $\phi_{R}$ directions is important, whereas parametric resonance and GW production are blind to $A\ll m_{\Phi}^{2}$, making radial coordinates more suitable.
+Interesting deviations from the $A=0$ limit are discussed in Section E of Supplemental Material.
+For $A\gtrsim 0.05m_{\Phi}^{2}$ the GW spectrum starts being sensitive to the exact value of $A$ as well as the initial angle $\theta_{\mathrm{in}}$.
+This provides intriguing possibilities for distinguishing between these parameters by dedicated analysis of the GW spectral shape.
+The growth of fluctuations and their backreaction breaks the homogeneous oscillation of $\Phi$, leading to inhomogeneities with an effective anisotropic stress tensor $\Pi_{ij}^{TT}=\nabla_{i}\phi^{a}\nabla_{j}\phi^{a}$ (where the repeated index a`a’ indicates the number of the scalar components and a summation is implied).
 The dynamics of the generated GWs are governed by the linearized equation for the transverse-traceless tensor perturbation sourced by the above anisotropic stress
-$\ddot{h}_{ij}+3H\dot{h}_{ij}-\frac{\nabla^{2}h_{ij}}{a^{2}}=\frac{2}{a^{2}M_{\mathrm{P}}^{2}}\Pi_{ij}^{TT},$ (14)
+$\ddot{h}_{ij}+3H\dot{h}_{ij}-\frac{\nabla^{2}h_{ij}}{a^{2}}=\frac{2}{a^{2}M_{\mathrm{P}}^{2}}\Pi_{ij}^{TT}\,.$ (15)
 The quantity of interest is the energy density power spectrum of the GWs normalized by the critical energy density of the Universe today.
@@
 (2024); Garcia-Bellido and Figueroa (2007); Garcia-Bellido et al.
-(2008):
+(2008)
 $\Omega_{\mathrm{G}W,0}h^{2}$ $=\frac{h^{2}}{\rho_{\mathrm{crit}}}\frac{d\rho_{\mathrm{G}W}}{d\ln k}\Bigg|_{t=t_{0}}=\frac{h^{2}}{\rho_{\mathrm{crit}}}\frac{d\rho_{\mathrm{G}W}}{d\ln k}\Bigg|_{t=t_{\mathrm{e}}}\frac{a_{\mathrm{e}}^{4}\rho_{\mathrm{e}}}{a_{0}^{4}\rho_{\mathrm{crit},0}}$
 $=\Omega_{\mathrm{rad},0}h^{2}\Omega_{\mathrm{GW},\mathrm{e}}\left(\frac{g_{\ast}}{g_{0}}\right)^{-1/3},$ (16)
-where the critical density of the Universe is $\rho_{\mathrm{crit}}=3M_{\mathrm{P}}^{2}H^{2}$ and $\Omega_{\mathrm{rad},0}h^{2}=h^{2}\rho_{\mathrm{rad},0}/\rho_{\mathrm{crit}}=4.3\times 10^{-5}$.
+where the critical density of the Universe is $\rho_{\mathrm{crit}}=3M_{\mathrm{P}}^{2}H^{2}$ and $\Omega_{\mathrm{rad},0}h^{2}=h^{2}\rho_{\mathrm{rad},0}/\rho_{\mathrm{crit}}\simeq 4.3\times 10^{-5}$.
 In particular, we employ the publicly available code Cosmolattice Figueroa et al.
@@
 (15) along with the equations for the scalar system in an expanding universe.
-The benchmark examples of observed spectra today from our model simulation are shown in Fig 1.
-For the initially $\Phi^{4}$ dominant system, as in our case, the field starts rolling/oscillating when $H\sim\sqrt{\lambda_{\Phi}}\phi_{\mathrm{in}}$.
-This combination of $\sqrt{\lambda_{\Phi}}\phi_{\mathrm{in}}$ also determines the frequency of the observed GWs, according to Eq.
-12.
-We consider the RD universe with the scalar field system contributing to a fraction of the total energy $\alpha$ and $\alpha=0.3$ is assumed in Fig 1.
-According to Eq.
-13, $\Omega^{\mathrm{peak}}_{\mathrm{GW}}\propto\alpha^{2}\propto\phi^{8}_{\mathrm{in}}$, and we found that a sizable $\alpha\sim O(0.1)$ which optimizes GW amplitude requires $\phi_{\mathrm{in}}\sim M_{\mathrm{P}}$.
+The spectra for three benchmark examples are shown in Fig 1.
+For a $\Phi^{4}$-dominated system, as in our case initially, the field starts rolling/oscillating when $H\sim\sqrt{\lambda_{\Phi}}\phi_{\mathrm{in}}$.
+The combination $\sqrt{\lambda_{\Phi}}\phi_{\mathrm{in}}$ also determines the frequency of the observed GWs.
+We consider the RD Universe with the scalar field system contributing a fraction of the total energy $\alpha$; Fig 1 corresponds to $\alpha=0.3$.
+For GWs generated by sources peaked around a characteristic scale $k_{c}$ in the momentum space Giblin and Thrane (2014), $\Omega^{\mathrm{peak}}_{\mathrm{GW}}\propto\alpha^{2}\propto\phi^{8}_{\mathrm{in}}$, and we found that a sizable $\alpha\sim O(0.1)$ which optimizes GW amplitude requires $\phi_{\mathrm{in}}\sim M_{\mathrm{P}}$ (see also Cui et al.
+(2024)).
+Furthermore, the observed frequency today is $f_{\mathrm{GW}}^{\mathrm{peak}}\simeq 2.7\times 10^{10}k_{c}/\sqrt{H_{c}M_{\mathrm{Pl}}}$ Hz, where $k_{c}\sim H\sim\sqrt{\lambda_{\Phi}}\phi_{\mathrm{in}}$ (see Section E of Supplemental Material for details), leading to $f_{\mathrm{GW}}^{\mathrm{peak}}\sim 10^{10}\lambda_{\Phi}^{1/4}$ for $\phi_{\mathrm{in}}\sim M_{\mathrm{Pl}}$.
 As we have seen from the discussion on the resonance structure of the model, the parameters $\mu_{\pm}$ (or simply $\mu$ when $A=0$) determine the nature and efficiency of resonance Greene et al.
 (1997).
-We take $\mu=10^{-5}$ as an example for stable resonance, which fixes $\lambda_{\Phi}\phi_{\mathrm{in}}^{2}$.
-with the optimal choice of $\phi_{\mathrm{in}}\sim M_{\mathrm{Pl}}$, we find $\lambda_{\Phi}\sim 10^{-30}$ for the mass range of interest.
+We found that stable resonance occurs for $\mu\equiv{m_{\Phi}}/({\sqrt{\lambda_{\Phi}}\phi_{\mathrm{in}}})\lesssim 10^{-5}$.
+We choose $\mu=10^{-5}$, which fixes the product $\lambda_{\Phi}\phi_{\mathrm{in}}^{2}$.
+With an optimal initial value of $\phi_{\mathrm{in}}\sim M_{\mathrm{Pl}}$, we find that $\lambda_{\Phi}\simeq 10^{-35},10^{-33},10^{-31}$ for $m_{\Phi}=0.1,1,10$ GeV; see Fig.
+1.
 As discussed earlier, $m_{\Phi}\gtrsim O(0.1)$ GeV is required for $\Phi$ to successfully transfer the asymmetry to SM baryons.
 Interestingly, in the viable range of $m_{\Phi}\sim O(0.1)-O(10)$ GeV, the resulting peak GW frequency lies well within the range of ET/CE and with sufficient amplitudes, making signal detection promising for the near future.
-A higher range of $m_{\Phi}$ is compatible with baryogenesis but requires futuristic higher frequency GW detectors Aggarwal et al.
+A higher range of $m_{\Phi}$ is compatible with baryogenesis but requires futuristic, higher-frequency GW detectors Aggarwal et al.
 (2021) to capture the peak frequency range of the signal.
@@
 (2007).
-Therefore, to extend the spectral prediction to lower frequencies, we apply the standard assumption that at low wave numbers, the GW spectrum grows as $k^{3}$, as derived from causality arguments Easther et al.
+Therefore, to extend the spectral prediction to lower frequencies, we apply the standard assumption that at low wave numbers, the GW spectrum scales as $k^{3}$, as derived from causality arguments Easther et al.
 (2007); Caprini et al.
+(2009); Cai et al.
 (2020); Hook et al.
@@
 1.
-Depending on $m_{\Phi}$, the low-$k$ tail may be detectable by DECIGO and BBO, providing the prospect for a correlated detection across multiple frequency bands, probing the peak and tail of the spectrum simultaneously.
+Depending on $m_{\Phi}$, the low-$k$ tail may be detectable by DECIGO and BBO, providing the prospect for correlated detection across multiple frequency bands, which simultaneously probes the peak and tail of the spectrum.
+Before concluding, let us note that the simulated signal shown in Fig.
+1 is ${\cal O}(100)$ times larger than the lowest point of the CE sensitivity curve.
+Since $\Omega_{GW}h^{2}\propto\alpha^{2}$ Giblin and Thrane (2014), we can reduce $\alpha$, the initial fraction of the energy density of the scalar field with respect to the total energy density in the Universe, down to $\alpha\sim O(1\%)$ and still yield a GW signal within reach of CE.
 Figure 1: We show the GW spectrum originating from the AD model considered here for three benchmark masses of $\Phi$.
-The potential parameters $\lambda_{\Phi}=10^{-30}$, $A=0$ are chosen to ensure efficient parametric resonance (see main text for details).
+The potential parameters $\lambda_{\Phi}\sim 10^{10}m_{\Phi}^{2}/M_{\mathrm{Pl}}^{2}$, $A=0$ are chosen to ensure efficient parametric resonance (see the main text for details).
 The fraction of initial scalar field energy density over the total energy density of the Universe is $\alpha=30\%$.
-The dashed lines at lower frequencies are extrapolated according to the causal super-horizon $k^{3}$ scaling.
-Finally, note that since $\Phi$ is coupled to SM fermions in order to transfer the asymmetry to baryons, the effect of parametric excitation of these fermionic states should be evaluated.
-Nevertheless, in the context of preheating, it was found that during scalar field oscillations, the energy transfer to bosons is typically faster and larger than to fermions due to Pauli blocking Greene and Kofman (1999); Giudice et al.
-(1999); Greene and Kofman (2000); Sorbo (2001); Adshead and Sfakianakis (2015).
-Therefore, we expect our resonance analysis to remain robust in the presence of fermion couplings.
+The dashed lines at lower frequencies are extrapolated according to the causal superhorizon $k^{3}$ scaling.
 Conclusion and Outlook.
 In this Letter, we first demonstrate that the Affleck-Dine mechanism for baryogenesis can be realized by a complex light scalar field $\Phi$ with a mass $m_{\Phi}$ well below the electroweak scale without invoking supersymmetry.
-In particular, we use a convex, renormalizable, polynomial potential (which does not lead to the formation of Q-balls generally present in SUSY AD models).
+In particular, we use a convex, renormalizable, polynomial potential (which does not lead to the formation of Q-balls, generally present in SUSY AD models).
 Such a scalar field acts as a spectator field during inflation while acquiring a large field value at the end of inflation.
@@
 Parametric resonance and fragmentation generally arise during the oscillation of the scalar field, naturally leading to GW production.
-We demonstrate that the theory may lead to successful baryogenesis while sourcing detectable GW signals with a peak frequency of O(10-100) Hz, within reach of experiments such as ET and CE.
+We demonstrate that the models considered can lead to successful baryogenesis while sourcing detectable GW signals with a peak frequency of O(10-100) Hz, within reach of experiments such as the ET and CE.
 Furthermore, the characteristic new physics scale, characterized by $m_{\Phi}$, intriguingly lies in the range of ${\cal O}(0.1\mathrm{-}10)$ GeV, while the transfer of the $\Phi$ asymmetry to the SM B- or L-asymmetry requires interactions between the new physics sector and the SM states.
@@
 The specifics of the complementary laboratory signal depend on the details of the asymmetry transfer mechanism.
-Examples include: searches for ${\cal O}(0.1\mathrm{-}10)$ GeV mass sterile neutrino with DUNE, SHiP, FASER, etc.
-and exotic multi-jet signals at the LHC (see Section .4 of the Supplemental Material).
+Examples include searches for ${\cal O}(0.1\mathrm{-}10)$ GeV mass sterile neutrino with DUNE, SHiP, FASER, etc., and exotic multijet signals at the LHC (see Supplemental Material, Section D).
 This work thus sheds new light on probing early Universe baryogenesis through its GW imprint, potentially complemented by associated laboratory probes, while presenting a well-motivated new physics source for SGWB.
@@
 The variance of the distribution grows as $\langle\phi^{2}\rangle=(H^{2}/4\pi^{2})N$ where $N$ is the $e$-folding number.
-We thus see that an equilibrium distribution with a Planckian standard deviation requires much more than the usual minimum amount of $60$ $e$-folds of inflation.
+We thus see that an equilibrium distribution with a Planckian standard deviation requires much more than the usual minimum amount of 6060 $e$-folds of inflation.
 This may be considered as a way to infer whether a prolonged inflationary phase took place in the early Universe.
@@
 (2024) for more details) applies to the absolute value of the complex scalar field $|\phi|$.
-The argument will enter its equilibrium distribution much sooner, and the distribution itself will be much simpler: uniform between 00 and $2\pi$.
+The argument will enter its equilibrium distribution much sooner, and the distribution itself will be much simpler: uniform between 0 and $2\pi$.
 The intricacies of multi-dimensional random walks and their applications to de Sitter fluctuations of fields with continuous symmetries (including complex scalars) can be found in Ref.
@@
 (2008).
-During inflation we can decompose the AD field $\Phi$ into a radial component $R$ and an angular component $\theta$.
+During inflation, we can decompose the AD field $\Phi$ into a radial component $R$ and an angular component $\theta$.
 We consider a typical value of the radial field $\bar{R}=\sqrt{\langle R^{2}\rangle}\sim M_{\mathrm{P}l}$ and thus
@@
 $n_{0}(t)=i\left(\Phi^{\dagger}\dot{\Phi}-\dot{\Phi}^{\dagger}\Phi\right)=\dot{\phi}_{\mathrm{R}}\phi_{{}_{\mathrm{I}}}-\phi_{{}_{\mathrm{R}}}\dot{\phi}_{\mathrm{I}}$ (S17)
-where we have used the subscript ‘00’ to denote the generated asymmetry in the $\Phi$ sector alone, without considering the decay to SM particles.
+where we have used the subscript ‘0’ to denote the generated asymmetry in the $\Phi$ sector alone, without considering the decay to SM particles.
 Substituting the approximate solutions for the background fields, the asymmetry becomes
@@
 $\tilde{n}_{{}_{C}}=\frac{\Gamma_{\Phi}\phi_{{}_{\ast}}^{2}\sin(2\theta)m_{{}_{\Phi}}^{2}}{2A}\left(1+\left(\frac{\Gamma_{\Phi}m_{{}_{\Phi}}}{2A}\right)^{2}\right)^{-1}.$ (S22)
-This expression has two limiting cases depending on whether the $\Phi$ lifetime is short ($\Gamma_{\Phi}\gg 2A/m_{{}_{\Phi}}$) or long ($\Gamma_{\Phi}\ll 2A/m_{{}_{\Phi}}$) compared to the period of oscillation of the asymmetry:
+This expression has two limiting cases, depending on whether the $\Phi$ lifetime is short ($\Gamma_{\Phi}\gg 2A/m_{{}_{\Phi}}$) or long ($\Gamma_{\Phi}\ll 2A/m_{{}_{\Phi}}$) compared to the period of oscillation of the asymmetry:
 $\tilde{n}_{{}_{C}}=\begin{cases}\frac{2A}{\Gamma_{\Phi}}\,\phi_{{}_{\ast}}^{2}\sin(2\theta)\quad&\Gamma_{\Phi}\gg 2A/m_{{}_{\Phi}}\\ \frac{\Gamma_{\Phi}m_{{}_{\Phi}}^{2}}{2A}\,\phi_{{}_{\ast}}^{2}\sin(2\theta)\quad&\Gamma_{\Phi}\ll 2A/m_{{}_{\Phi}}\end{cases}\,.$ (S23)
@@
 $\frac{n_{B}}{s}\Bigg|_{\mathrm{obs}}=0.861\pm 0.005\times 10^{-10}.$
-To this end, we multiply the comoving transferred asymmetry with $(a_{\ast}/a)^{3}$ to compute the total asymmetry
+To this end, we multiply the comoving transferred asymmetry by $(a_{\ast}/a)^{3}$ to compute the total asymmetry
 $\tilde{n}=\left(\frac{a_{\ast}}{a}\right)^{3}\tilde{n}_{{}_{C}}$ (S24)
-For the RD universe, as in our case, the equation of state is $w=1/3$ leading to
+For the RD Universe, as in our case, the equation of state is $w=1/3$, leading to
 $\left(\frac{a_{\ast}}{a}\right)^{3}=\left(\frac{H(t)}{H_{\ast}}\right)^{3/2}$ (S25)
 Dividing $\tilde{n}$ by the entropy density $s=4k_{T}^{2}T^{3}$, with $k_{T}=\sqrt{\pi^{2}g(T)/90}$, we arrive at the baryon-to-photon ratio
-$\frac{n_{B}}{s}=\begin{cases}\left(\frac{4\alpha^{3}}{k_{T_{d}}^{6}\lambda_{{}_{\Phi}}}\right)^{1/4}\left(\frac{A}{m_{{}_{\Phi}}^{2}}\right)\left(\frac{m_{{}_{\Phi}}M_{\mathrm{P}}}{T_{d}^{2}}\right)\sin(2\theta)\quad&\Gamma_{\Phi}\gg 2A/m_{{}_{\Phi}}\\ \left(\frac{\alpha^{3}k_{T_{d}}^{2}}{64\lambda_{{}_{\Phi}}}\right)^{1/4}\left(\frac{m_{{}_{\Phi}}^{2}}{A}\right)\left(\frac{T_{d}^{2}}{m_{{}_{\Phi}}M_{\mathrm{P}}}\right)\sin(2\theta)\quad&\Gamma_{\Phi}\ll 2A/m_{{}_{\Phi}}\,,\end{cases}$ (S26)
+[…]
 where the $\Phi$ decay completes when $\Gamma_{\Phi}=H(T_{d})=k_{T_{d}}T_{d}^{2}/M_{\mathrm{P}}$ and the Hubble scale at $t_{*}$ is $H_{\ast}\simeq\frac{3}{4}{m_{{}_{\Phi}}^{2}}/{\lambda_{{}_{\Phi}}}$.
 We can also elucidate the dependence on $\phi_{\mathrm{in}}$, by rewriting the above expression as
-$\frac{n_{B}}{s}=\begin{cases}\mu\left(\frac{4\lambda_{{}_{\Phi}}\alpha^{3}}{k_{T_{d}}^{6}}\right)^{1/4}\left(\frac{A}{m_{{}_{\Phi}}^{2}}\right)\left(\frac{\phi_{\mathrm{in}}M_{\mathrm{P}}}{T_{d}^{2}}\right)\sin(2\theta)\quad&\Gamma_{\Phi}\gg 2A/m_{{}_{\Phi}}\\ \mu\left(\frac{\alpha^{3}k_{T_{d}}^{2}}{64\lambda_{{}_{\Phi}}^{3}}\right)^{1/4}\left(\frac{m_{{}_{\Phi}}^{2}}{A}\right)\left(\frac{T_{d}^{2}}{\phi_{\mathrm{in}}M_{\mathrm{P}}}\right)\sin(2\theta)\quad&\Gamma_{\Phi}\ll 2A/m_{{}_{\Phi}}\,,\end{cases}$ (S27)
+[…]
 where for sufficient resonance, we require $\mu^{2}\equiv m_{{}_{\Phi}}^{2}/(\lambda_{{}_{\Phi}}\phi_{\mathrm{in}}^{2})\lesssim 10^{-5}$.
-In the case where the asymmetry in the $\Phi$ field is first transferred to a SM lepton asymmetry $n_{L}$, which is later converted to a baryon asymmetry $n_{B}$ via the sphaleron process, the above result gets multiplied by a factor of $28/79$ since Harvey and Turner (1990); Asaka et al.
-(2002); Garcia-Bellido et al.
-(1999)
-$\frac{n_{B}}{s}=-\frac{28}{79}\frac{n_{L}}{s}\,.$ (S28)
-Figure S1 shows the resulting asymmetry as a function of the temperature of the universe at the time of complete $\Phi$ decay for different values of the potential parameters; $m_{\Phi}$ and $A$.
+When written in terms of the dimensionless parameters, $\epsilon_{\Phi}=A/m_{\Phi}^{2}$ and $\gamma_{\Phi}=\Gamma_{\Phi}/m_{\Phi}$, we find the following familiar expression:
+$\frac{n_{B}}{s}=\begin{cases}\left(\frac{4\alpha^{3}}{\lambda_{\Phi}k_{T_{d}}^{6}}\right)^{1/4}\epsilon_{\Phi}\frac{m_{\Phi}M_{\mathrm{Pl}}}{T_{d}^{2}}\sin(2\theta);&\gamma_{\Phi}\gg 2\epsilon_{\Phi}\\ \left(\frac{\alpha^{3}k_{T_{d}}^{2}}{64\lambda_{\Phi}}\right)^{1/4}\frac{1}{\epsilon_{\Phi}}\frac{T_{d}^{2}}{m_{\Phi}M_{\mathrm{Pl}}}\sin(2\theta);&\gamma_{\Phi}\ll 2\epsilon_{\Phi}\end{cases}$ (S28)
+Figure S1 shows the resulting asymmetry, following Eq.
+(S22), as a function of the temperature of the Universe at the time of complete $\Phi$ decay for different values of the potential parameters; $m_{\Phi}$ and $A$.
 The quartic term is chosen to keep $m_{\Phi}^{2}/\lambda_{\Phi}\phi_{\mathrm{in}}^{2}=10^{-5}$ for sufficient resonance.
 We see the strong dependence of $n_{B}/s$ on $A$ and $T_{d}$, as expected.
-This allows the model to be viable, even in cases where significant washout occurs, by increasing the initially produced asymmetry through proper parameter choices.
-Finally, let us reiterate one assumption that was made in the above analytic results for the generated asymmetry.
+This enables the model to remain viable, even in cases where significant washout occurs, by increasing the initially produced asymmetry through appropriate parameter choices.
+Finally, let us reiterate one assumption made in the above analytic results regarding the generated asymmetry.
 As mentioned in the main text, our analysis assumes that the motion of $\Phi$ is radial at early times, for $\lambda_{\Phi}|\Phi|^{4}\gg m_{\Phi}^{2}|\Phi|^{2}$.
-After that, the potential is dominated by the quadratic terms, which in our case includes the bare mass and the asymmetry $A(\Phi^{2}+\Phi_{*}^{2})$.
+After that, the potential is dominated by the quadratic terms, which in our case include the bare mass and the asymmetry $A(\Phi^{2}+\Phi_{*}^{2})$.
 By simply equating the quartic and bare mass terms, we arrive at the time of transition between the two regimes of $t_{*}$ when $|\Phi(t_{*})|=m_{\Phi}/\sqrt{\lambda_{\Phi}}$.
-Demanding that at early times (before $t_{*}$) the asymmetry term can be neglected leads to $\epsilon_{\Phi}=A/m_{\phi}^{2}<(\mu^{2}/15\alpha)(\phi_{\mathrm{in}}^{2}/M_{\mathrm{Pl}}^{2})$ where $\mu^{2}\equiv m_{\Phi}^{2}/(\lambda_{\Phi}\phi_{\mathrm{in}}^{2})\lesssim 10^{-5}$.
+Demanding that at early times (before $t_{*}$) the asymmetry term can be neglected leads to $\epsilon_{\Phi}<(\mu^{2}/15\alpha)(\phi_{\mathrm{in}}^{2}/M_{\mathrm{Pl}}^{2})$ where $\mu^{2}\equiv m_{\Phi}^{2}/(\lambda_{\Phi}\phi_{\mathrm{in}}^{2})\lesssim 10^{-5}$.
 For $\phi_{\mathrm{in}}\sim M_{\mathrm{Pl}}$ the above inequality can be written as $\epsilon_{\Phi}\lesssim 10^{-5}$.
-If we push $\epsilon_{\Phi}$ above this value, the model is not invalided, but the asymmetry would get an extra suppression factor since $\sin\theta<\sin\theta_{\mathrm{in}}$.
+If we push $\epsilon_{\Phi}$ above this value, the model is not invalidated, but the asymmetry would get an extra suppression factor since $\sin\theta<\sin\theta_{\mathrm{in}}$.
 The amount of suppression depends on the amount of non-radial motion for $t<t_{*}$.
 That being said, the computed $n_{B}/s$ can be many orders of magnitude above the measured value of $10^{-10}$, meaning that a suppression of $\sin\theta$ is not “fatal” for the model, even for $\epsilon_{\Phi}>10^{-5}$.
-We do not perform a detailed analysis for this regime, which simply involves computing the angular motion of $\Phi$ from the first instant when it starts rolling for a given set of potential parameters.
+We do not perform a detailed analysis for this regime, which involves simply computing the angular motion of $\Phi$ from the first instant when it starts rolling, given a specific set of potential parameters.
 As an example, for the benchmark values of $\lambda_{\Phi}\sim 10^{-30}$ and $m_{\Phi}=10\,\mathrm{GeV}$, we must have $\epsilon_{\Phi}\lesssim 10^{-6}$ for the validity of our estimates in Eq.
+(S27).
+The temperature at the decay can vary between $T_{d}^{\mathrm{max}}\sim 10^{8}\,\mathrm{GeV}$ for instantaneous decay and $T_{d}^{\mathrm{min}}=10\mathrm{MeV}$, if the decay occurs right before BBN Kawasaki et al.
+(1999, 2000); Steigman (2007); Hasegawa et al.
 (2019).
-The temperature at the decay can vary between $T_{d}^{\mathrm{max}}\sim 10^{8}\,\mathrm{GeV}$ for instantaneous decay and $T_{d}^{\mathrm{min}}=10\mathrm{MeV}$, if the decay occurs right before BBN.
 Therefore, to obtain the correct ratio of $n_{B}/s$, we require $\epsilon_{\Phi}=10^{-38}(T_{d}/\mathrm{GeV})^{2}$ when $\gamma_{\Phi}\ll\epsilon_{\Phi}$.
-Figure S1: The maximum baryon-to-photon ratio $\left.{n_{B}/s}\right|_{\mathrm{max}}={n_{\Phi}/s}$ (neglecting possible washout and the sphaleron factor) for three different masses of the scalar field as in Fig (1) as a function of the temperature $T_{d}$ of the $\Phi$ decay.
+Figure S1: The maximum baryon-to-photon ratio $\left.{n_{B}/s}\right|_{\mathrm{max}}={n_{\Phi}/s}$ (neglecting possible washout and the sphaleron factor) for three different masses of the scalar field, as in Fig.
+(1) of the main text, as a function of the temperature $T_{d}$ of the $\Phi$ decay.
 The solid and dashed lines correspond to $\epsilon_{\Phi}=10^{-20}$ and $\epsilon_{\Phi}=10^{-6}$ respectively.
@@
 .4 Transfer of the $\Phi$ asymmetry to the SM baryon asymmetry
-In this section we elaborate the example models realizing the transfer of $\Phi$ asymmetry to the SM baryon asymmetry and the related constraints.
+In this section, we elaborate on the example models realizing the transfer of $\Phi$ asymmetry to the SM baryon asymmetry and the related constraints.
 (A) $\Phi$ decay to leptons (leptogenesis).
@@
 (2022); Shuve and Yavin (2014); Hambye and Teresi (2016).
-As seen from our GW results, $m_{\Phi}\lesssim O(10)$ GeV is preferred in order to yield detectable GW frequency below $\sim 100$ Hz.
+As seen from our GW results, $m_{\Phi}\lesssim O(10)$ GeV is preferred in order to yield a detectable GW frequency below $\sim 100$ Hz.
 This implies that light $N$ of mass $\lesssim O(1-10)$ GeV.
 In this mass range, conventional $N$ decays to SM leptons via a 3-body or loop-suppressed process, which typically occurs after electroweak sphaleron already turns off, albeit before BBN.
-The washout effect from inverse decay or back-scattering is also significant in this low mass range.
+The washout effect from inverse decay or backscattering is also significant in this low-mass range.
 However, as found in recent leptogenesis literature, asymmetry transfer from $N$ to SM leptons can be realized while sphaleron is still active by freeze-out (generalized out-of-equilibrium decay or scattering) or freeze-in (via $N$-$\nu_{\mathrm{S}M}$ mixing/oscillation) mechanisms, for $m_{N}$ as low as O(100) MeV.
@@
 Even in the presence of large washout suppression, as shown in the main text, our Affleck-Dine model can initially generate a larger $\Phi$ asymmetry, which compensates for the washout to eventually result in the observed baryon asymmetry.
-Unlike with the leptogeneis case, here, the $NHL$ interaction only serves the purpose of transferring the $\Phi$ ($N$) asymmetry to SM leptons instead of generating the asymmetry via CP violation in the SM neutrino sector.
-Therefore, the details of the asymmetry evolution in our case may differ from its leptogenesis counterparts and are highly model-dependent, which may be elucidated by dedicated studies beyond the scope of this work.
+Unlike with the leptogenesis case, here, the $NHL$ interaction only serves the purpose of transferring the $\Phi$ (NN) asymmetry to SM leptons instead of generating the asymmetry via CP violation in the SM neutrino sector.
+Therefore, the details of the asymmetry evolution in our case may differ from those of its leptogenesis counterparts and are highly model-dependent, which can be elucidated by dedicated studies beyond the scope of this work.
 Furthermore, an alternative to the Majorana $N$ as shown in Eq.
-S29 is to introduce quasi-Dirac $N$ where the washout effect can be more suppressed for GeV mass range (also see Cline et al.
+S29 is to introduce quasi-Dirac $N$, where the washout effect can be more suppressed for the GeV mass range (also see Cline et al.
 (2020b) for this possibility and the related phenomenological implications).
@@
 This leads to the constraint $y_{\Phi}\gtrsim 10^{-7}\sqrt{\frac{\mathrm{GeV}}{m_{\Phi}}}$, which can be easily satisfied.
-Laboratory signatures of the case of $\Phi$ decaying to leptons generally relate to sterile neutrino searches in the mass range of $O(0.1-10)$ GeV, which are being conducted at a range of experiments such as DUNE, SHiP, FASER et al.
+Laboratory signatures of the case of $\Phi$ decaying to leptons generally relate to sterile neutrino searches in the mass range of O(0.110)O(0.1-10) GeV, which are being conducted at a range of experiments such as DUNE, SHiP, FASER et al.
 ste (2022); Bolton et al.
@@
 $\mathcal{L}\supset y_{\chi}\phi_{\chi}\chi\chi+g_{i}\chi\bar{u}_{i}\varphi+\lambda_{ij}\varphi d_{i}d_{j}+{\mathrm{h}.c.}$ (S30)
-where $\chi$ is a singlet Majorana fermion, $\varphi$ is an up-type diquark scalar, and the flavor indices $i,~{}j=1,2,3$ in down-type quark combination $d_{i}d_{j}$ should be antisymmetric.
-This type of model finds its close analogy in supersymmetric theories, in particular, $\chi$ may relate to neutralino or singlino Cui and Sundrum (2013), $\varphi$ an up-type squark, $\lambda_{ij}$ as UDD type of R-parity violating (RPV) coupling.
-Therefore, existing searches for RPV SUSY may lead to constraints on masses and couplings that we consider here.
-The LHC constraints on the colored particle $\phi$ depend on whether it resembles the first two generations of squark or stop, and depends on whether RPV decay $\phi\rightarrow dd$ or R-parity conserving decay $\phi\rightarrow u\chi$ dominates (assuming $\chi$ reveals itself as MET) Karmakar (2023); Sirunyan et al.
+where $\chi$ is a singlet Majorana fermion, $\varphi$ is an up-type diquark scalar, and the flavor indices i,j=1,2,3i,~j=1,2,3 in down-type quark combination $d_{i}d_{j}$ should be antisymmetric.
+This type of model finds its close analogy in supersymmetric theories.
+In particular, $\chi$ may relate to a neutralino or singlino Cui and Sundrum (2013), $\varphi$ to an up-type squark, and $\lambda_{ij}$ to a UDD-type R-parity-violating (RPV) coupling.
+Therefore, existing searches for RPV SUSY may lead to constraints on masses and couplings that we consider in this context.
+The LHC constraints on the colored particle $\phi$ depend on whether it resembles the first two generations of squark or stop, and depend on whether RPV decay $\phi\rightarrow dd$ or R-parity conserving decay $\phi\rightarrow u\chi$ dominates (assuming $\chi$ reveals itself as MET) Karmakar (2023); Sirunyan et al.
 (2018, 2021); Aad et al.
@@
 In any case, generally, $m_{\varphi}\gtrsim$1 TeV is the ballpark of the current LHC constraints, which we will use for our estimates.
-This constraint determines that in the parameter range of our interest where $m_{\chi}\sim O(1-10)$ GeV, $\chi$ decays proceeds with 3-body channel $\chi\rightarrow udd$.
+This constraint determines that in the parameter range of our interest, where $m_{\chi}\sim O(1-10)$ GeV, $\chi$ decays proceed with the 3-body channel $\chi\rightarrow udd$.
 For simplicity, we assume $g_{i}\sim 1$.
@@
 (2013).
-Due to the antisymmetry in $i,~{}j$, $n$-$\bar{n}$ constraint is rather weak, while di-nucleon decay $pp\rightarrow K^{+}K^{+}$ provides the strongest bound Goity and Sher (1995); Barbier et al.
+Due to the antisymmetry in i,ji,~j, $n¯\bar{n}$ constraint is rather weak, while di-nucleon decay $pp\rightarrow K^{+}K^{+}$ provides the strongest bound Goity and Sher (1995); Barbier et al.
 (2005) of $\lambda_{12}\lesssim 10^{-6}$, for $m_{\varphi}\sim$ TeV, $m_{\chi}\sim$ O(GeV).
-However, the bound is relaxed for heavy flavor quarks, e.g.
-$\lambda_{23}\lesssim 1$.
-The kinematic condition of $\chi$ decaying to baryons leads to lower limit on masses: $m_{\Phi}>2$ GeV, $m_{\chi}>1$ GeV.
+However, the bound is relaxed for heavy flavor quarks, e.g., $\lambda_{23}\lesssim 1$.
+The kinematic condition of $\chi$ decaying to baryons leads to a lower limit on masses: $m_{\Phi}>2$ GeV, $m_{\chi}>1$ GeV.
 The decays of both $\Phi\rightarrow\chi\chi$ and $\chi\rightarrow udd$ are required to occur before BBN, i.e., $\Gamma_{\Phi\rightarrow\chi\chi}\sim\frac{1}{16\pi}y_{\chi}^{2}m_{\Phi}>H(T_{\mathrm{BBN}})$, and $\Gamma_{\chi\rightarrow udd}>H(T_{\mathrm{BBN}})$.
@@
 If $\lambda_{12}$ dominates $\chi$ decay, this would contradict the aforementioned di-nucleon decay bound.
-Therefore, we choose to consider the pattern of heavy flavor domination, which is well motivated and has been widely considered in SUSY context Brust et al.
+Therefore, we choose to consider the pattern of heavy flavor domination, which is well-motivated and has been widely studied in the SUSY context Brust et al.
 (2012a, b); Cui et al.
@@
 (2018).
-.5 Asymmetry dependence of GW production
+.5 Parametric resonance and asymmetry dependence of GW production
 It is worth exploring in more detail the connection between the asymmetry of the potential and the shape and amplitude of the resulting GW spectrum.
+The growth of fluctuations and their backreaction breaks the homogeneous oscillation of $\Phi$, leading to scalar field inhomogeneities with an effective anisotropic stress tensor $\Pi_{ij}^{TT}=\nabla_{i}\phi^{b}\nabla_{j}\phi^{b}$ (where the repeated index b``b" indicates the number of the scalar components and a summation is implied).
+These time-dependent field inhomogeneities source (sub-Hubble) gravitational waves whose frequency depends on the characteristic energy scale involved $H_{c}$.
+The peak amplitude and frequency of GW spectra generated from sources peaked at some characteristic scale $k_{c}$ in momentum space can be estimated as Giblin and Thrane (2014)
+$f_{\mathrm{GW}}^{\mathrm{peak}}=$ $2.7\times 10^{10}\sqrt{\frac{H_{c}}{M_{\mathrm{Pl}}}}\frac{k_{c}}{H_{c}}\,{\mathrm{Hz}}\,$, (S31)
+$\Omega_{\mathrm{GW}}^{\mathrm{peak}}=$ $2.3\times 10^{-4}\,\alpha^{2}\,\beta\,w\left(\frac{k_{c}}{\sigma}\right)\left(\frac{H_{c}}{k_{c}}\right)^{2}\,$, (S32)
+where $\alpha$ is a measure of the fraction of the energy in the GW source – here the scalar field inhomogeneities – relative to the total energy density of the Universe, and $\beta$ encodes the degree of anisotropy of the source, which must be determined from simulations.
+In typical scenarios $\beta=\mathcal{O}(0.01-0.1)$.
+The equation of the state of the Universe at the time of the production of GWs for a radiation-dominated Universe is $w=1/3$ and $\sigma$ is the width of the source in momentum space, which can be taken as $\sigma\sim k_{c}$ for peaked sources (see Ref.
+Giblin and Thrane (2014) for the derivation and e.g.
+Refs.
+Cui and Sfakianakis (2023); Cui et al.
+(2024) for further applications of Eqs (S31), (S32)).
 Let us start from the symmetric case $A=0$, where the potential becomes
@@
 Since the potential is radially symmetric, we can redefine the fields as $\phi_{\parallel}=\phi_{R}\cos\theta-\phi_{I}\sin\theta$ and $\phi_{\perp}=\phi_{R}\sin\theta+\phi_{I}\cos\theta$, where at the background level $\bar{\phi}_{\perp}=0$.
-This essentially makes the analysis follow our earlier work Cui et al.
-(2024), where we considered (among others) systems with two scalar fields with a quartic coupling, where one field is initially displaced from its minimum, and the other is (classically) at zero.
-This specific example is a combination of Models A and D of Ref.
+This essentially follows our earlier work Cui et al.
+(2024), where we considered (among others) systems with two scalar fields featuring a quartic coupling, with one field initially displaced from its minimum and the other (classically) at zero.
+This specific example combines Models A and D from Ref.
 Cui et al.
@@
 The equations of motion for the fluctuations in the $\phi_{\parallel}-\phi_{\perp}$ basis are
-Figure S2: The spectra of the scalar field fluctuations along (left) and perpendicular (right) to the direction of the background trajectory in the case of a rotationally symmetric potential with $A=0$, $\lambda_{\Phi}=5\times 10^{-35}$ and $m_{\Phi}=0.1$ GeV.
+Figure S2: The spectra of the scalar field fluctuations along (left) and perpendicular (right) to the direction of the background trajectory in the case when the potential has a rotation symmetry with $A=0$, $\lambda_{\Phi}=5\times 10^{-35}$ and $m_{\Phi}=0.1$ GeV.
 $\ddot{\delta\phi_{\parallel}}+3H\dot{\delta\phi_{\parallel}}+\left({k^{2}\over a^{2}}+m_{\Phi}^{2}+3\lambda_{\Phi}\bar{\phi}^{2}\right)\delta\phi_{\parallel}=0$ (S34)
 $\ddot{\delta\phi_{\perp}}+3H\dot{\delta\phi_{\perp}}+\left({k^{2}\over a^{2}}+m_{\Phi}^{2}+\lambda_{\Phi}\bar{\phi}^{2}\right)\delta\phi_{\perp}=0$ (S35)
-In principle, we can see two sources of parametric resonance: self-resonance of the ϕsubscriptitalic-$\phi_{\parallel}$ field due to the quartic term $3\lambda_{\Phi}\bar{\phi}^{2}\delta\phi_{\parallel}$ and resonant amplification of the ϕsubscriptitalic-$\phi_{\perp}$ field through the coupling term $\lambda_{\Phi}\bar{\phi}^{2}\delta\phi_{\perp}$.
+In principle, we can see two sources of parametric resonance: self-resonance of the $\phi_{\parallel}$ field due to the quartic term $3\lambda_{\Phi}\bar{\phi}^{2}\delta\phi_{\parallel}$ and resonant amplification of the $\phi_{\perp}$ field through the coupling term $\lambda_{\Phi}\bar{\phi}^{2}\delta\phi_{\perp}$.
 The background field trajectory is (by construction) along $\bar{\phi}_{\perp}=0$ and thus a straight line through the origin on the $\phi_{R}-\phi_{I}$ plane.
@@
 (1997), where $q$ is defined as the coefficient before $\lambda_{\Phi}\bar{\phi}^{2}$ in Eqs.
-S33.
+S35.
 Figure S2 shows the spectra of $\delta\phi_{\parallel},\delta\phi_{\perp}$ for different times, where we see that the $\delta\phi_{\perp}$ spectrum dominates, as expected from the standard analysis of preheating of quartic fields Greene et al.
@@
 In the $\phi_{R}-\phi_{I}$ basis, the spectra will be identical for $\theta=\pi/4$, since each spectrum has an equal contribution from $\delta\phi_{\parallel}$ and $\delta\phi_{\perp}$.
-If the initial angle between the real and imaginary parts of $\Phi$ changes, the spectra of $\delta\phi_{\parallel}$ and $\delta\phi_{\perp}$ remain unchanged, whereas the spectra of $\phi_{R}$ and $\phi_{I}$ change, since the inner product between the basis vectors of the two bases change with the angle $\theta$.
+If the initial angle between the real and imaginary parts of $\Phi$ changes, the spectra of $\delta\phi_{\parallel}$ and $\delta\phi_{\perp}$ remain unchanged, whereas the spectra of $\phi_{R}$ and $\phi_{I}$ change, since the inner product between the basis vectors of the two bases changes with the angle $\theta$.
 That being said, the resulting GW spectrum is unchanged, as expected, due to the rotational symmetry of the potential.
@@
 We parametrize the effect by $\epsilon_{\Phi}=A/m_{\Phi}^{2}$.
-As expected, small values of the angle do not significantly affect the background trajectory and the resonance structure.
+Let us first discuss the background trajectory.
+For $\epsilon_{\Phi}=0$ it is a straight line segment through the origin, with a length of ${\cal O}(\phi_{\mathrm{in}})$ which will shrink in time due to red-shifting.
+For $0<\epsilon_{\Phi}\ll 1$, the trajectory during each period will resemble an ellipse.
+Using Eqs.
+(S15)-(S16)) and Taylor-expanding for $\epsilon_{\Phi}\ll 1$, we can see that the minor axis of the (approximate) ellipse scales as ${\cal O}(\phi_{\mathrm{in}}\cdot A/m_{\phi}^{2})$, whereas the major axis is simply ${\cal O}(\phi_{\mathrm{in}})$.
+The ellipticity is then $\sim(A/m_{\phi}^{2})\ll 1$ for all cases of interest shown in the main text.
+For small values of $\epsilon_{\Phi}$, the background trajectory will be an ellipse with a very small ellipticity, and it will therefore behave almost like a straight line segment.
+We thus expect the parametric resonance results to be unaffected by the exact value of $\epsilon_{\Phi}\ll 1$.
 As seen in Fig.
-S3, for $\epsilon_{\phi}\leq 0.005$, the GW spectrum is independent of the initial angle $\theta_{\mathrm{in}}$ and actually indistinguishable from the symmetric case with $\epsilon_{\Phi}=0$.
+S3, for $\epsilon_{\phi}\leq 0.005$, the GW spectrum is independent of the initial angle $\theta_{\mathrm{in}}$ and indistinguishable from the symmetric case with $\epsilon_{\Phi}=0$.
 This means that for the asymmetry values that we use for our benchmark examples $\epsilon_{\Phi}<10^{-5}$, the resonance is “blind” with respect to the exact value of asymmetry in the potential.
-Figure S3: The resulting GW spectra for different values of the asymmetry parameter $A$ (color-coded according to the legend) and initial angle $\theta\equiv\arctan(\phi_{I}/\phi_{R})|_{\mathrm{init}}=\pi/6,\pi/4,\pi/3$ (solid, dashed and dotted respectively).
+Figure S3: The resulting GW spectra for different values of the asymmetry parameter AA (color-coded according to the legend) and initial angle $\theta\equiv\arctan(\phi_{I}/\phi_{R})|_{\mathrm{init}}=\pi/6,\pi/4,\pi/3$ (solid, dashed and dotted respectively).
 That being said, it is worth pushing the asymmetry value higher in order to cover the entire possible parameter space.
-For $0.05\lesssim\epsilon_{\Phi}\lesssim 0.15$ we see an enhancement of the GW spectrum at large values of the frequency $f\gtrsim 100$ Hz, while the independence of $\theta_{\mathrm{in}}$ still holds.
+For $0.05\lesssim\epsilon_{\Phi}\lesssim 0.15$, we see an enhancement of the GW spectrum at large values of the frequency $f\gtrsim 100$ Hz, while the independence of $\theta_{\mathrm{in}}$ still holds.
 This is an intriguing observation, as it provides an additional motivation for GW detectors that can probe frequencies around $1$ kHz with a sensitivity similar to CE and ET Zhang et al.
@@
 (2023).
-For $\epsilon_{\Phi}\simeq 0.25$ the UV enhancement of the GW spectrum remains, but we also see a difference between the initial angle choices between the real and imaginary parts of $\Phi$.
+For $\epsilon_{\Phi}\simeq 0.25$, the UV enhancement of the GW spectrum remains, but we also see a difference between the initial angle choices between the real and imaginary parts of $\Phi$.
 This means that a possible detection of a spectrum with such a UV enhancement would suffer from a partial degeneracy between $\epsilon_{\phi}$ and $\theta_{\mathrm{in}}$.
@@
 Finally, for $\epsilon_{\Phi}\gtrsim 0.3$, parametric resonance is severely suppressed, rendering the GW signals below the detection threshold.
+Before concluding, we must note that since $\Phi$ is coupled to SM fermions in order to transfer the asymmetry to baryons, the effect of parametric excitation of these fermionic states should be evaluated.
+Nevertheless, in the context of preheating, it was found that during scalar field oscillations, the energy transfer to bosons is typically faster and larger than to fermions due to Pauli blocking Greene and Kofman (1999); Giudice et al.
+(1999); Greene and Kofman (2000); Sorbo (2001); Adshead and Sfakianakis (2015).
+Therefore, we expect our resonance analysis to remain robust in the presence of fermion couplings.